### v1

# ../Bibtex/sources.bib

Abstract.
We derive explicit equations for the Oguiso-Yu automorphism of minimum topological entropy on a complex Enriques surface. The approach is computer aided and makes use of elliptic fibrations.
1. Introduction
Let $X$ be a smooth, connected, compact complex Kähler surface and $f\in\mathrm{Aut}(X)$ be an automorphism. The topological entropy $h(f)$ of $f$ measures the dynamical complexity of $f$. By theorems of Gromov and Yomdin [Yomdin87, gromov90] it is the logarithm
$h(f)=\log d_{1}(f)$
of the first dynamical degree $d_{1}(f)$. The first dynamical degree is a bimeromorphic invariant and agrees with the spectral radius $\lambda(f)$ of the pullback $f^{*}\in\mathrm{GL}(H^{2}(X,\mathbb{C}))$. It is either equal to one or a so-called Salem number. The latter is a real algebraic integer $\lambda>1$ which is Galois conjugate to $\lambda^{-1}$ and such that all other conjugates lie on the unit circle.
Conjecturally, the smallest Salem number is Lehmer’s number $\lambda_{10}\approx 1.17628$. The conjecture is true for the subset of Salem numbers coming from surface automorphisms: By a result of McMullen, the set of entropies of surface automorphisms has a spectral gap [McMullen2007, Theorem 1.2], i.e.
$h(f)=0\quad\mbox{or}\quad h(f)\geq\log\lambda_{10}.$
Moreover, there exists a rational surface and a projective K3 surface with an automorphism of minimal entropy given by Lehmer’s number $\lambda_{10}$ [McMullen2016].
The proof of existence for the K3 case works by constructing an integral Hodge structure $H$ of weight two and an effective Hodge isometry $f^{\prime}\in\mathrm{GL}(H)$. Then the surjectivity of the period map and the strong Torelli theorem imply that there exist a K3 surface $X$, a marking $\eta\colon H^{2}(X,\mathbb{Z})\to H$ and an automorphism $f\in\mathrm{Aut}(X)$ realising this data, i.e. with $\eta\circ f^{*}=f^{\prime}\circ\eta$. Being of transcendental nature, the Torelli theorem is a mere existence result. It does not provide explicit equations. In [BrandhorstElkies23] the first named author and Elkies give explicit equations for an automorphism $f$ with minimal entropy of a K3 surface $X$. The approach there was to make use of the different elliptic fibrations of $X$.
Cantat [Cantat1999, Proposition 1] observed that if $d_{1}(f)>1$, then $X$ is bimeromorphic to either a rational surface, a 2-torus, a K3 surface or an Enriques surface. For $2$-tori the set of entropies is known thanks to results of Reschke [Reschke2012, Reschke2017]. Oguiso and Yu [OguisoYu20] show that for an Enriques surface $Y$ and an automorphism $\widetilde{f}\in\mathrm{Aut}(Y)$, its entropy $h(\widetilde{f})$ is either zero or at least $\log\tau_{8}\approx 1.58234$ where $\tau_{8}$ is a root of the Salem polynomial
$01-x^{2}-2x^{3}-x^{4}+x^{6}=$ 0 .
Moreover, they prove the existence of a (complex) Enriques surface $Y_{OY}$ and an automorphism $\widetilde{f}_{OY}\in\mathrm{Aut}(Y_{OY})$ with entropy $h(\widetilde{f}_{OY})=\log\tau_{8}$ by working on the Hodge-theoretic side. Again, this is an abstract existence result, since the proof follows McMullen’s strategy invokes the Torelli Theorem for Enriques surfaces. Enriques surfaces are projective; therefore Oguiso and Yu ask in [OguisoYu20, Question 1.3] for explicit equations of $Y_{OY}$ and $\widetilde{f}_{OY}$. The purpose of this work is to answer their question and to popularise the methods developed along the way.
The code for reproduction of the results in this article is available on Zenodo [BrandhorstZach25code]. We will in the following refer to the files and their names as uploaded there.
Recall that the universal cover of a complex Enriques surface $Y$ is a K3 surface $X$ and that the projection $X\to Y$ is of degree $2$. Let $\iota\in\mathrm{Aut}(X)$ denote the covering involution (in other words, the non-trivial deck transformation). It is fixed-point free. Conversely, a fixed-point free involution on a K3 surface $X$ is called an Enriques involution, since the quotient $X/\iota$ is an Enriques surface. We denote by $X_{OY}$ the universal cover of $Y_{OY}$.
Theorem 1.1.
The Oguiso-Yu K3 surface $X_{OY}$ is the elliptic K3 surface whose weierstrass equation is
$y^{2}=x^{3}+(1-t^{8})x.$
Translation by the $2$-torsion section $(x,y)=(0,0)$ composed with $(x,y,t)\mapsto(x,y,-t)$ defines an Enriques involution $\iota_{OY}$. Then $Y_{OY}=X_{OY}/\iota_{OY}$ is the Oguiso-Yu Enriques surface.
It admits an automorphism $\widetilde{f}_{OY}\in\mathrm{Aut}(Y_{OY})$ of entropy $h(\widetilde{f}_{OY})=\log\tau_{8}$ which is defined over the number field $K=\mathbb{Q}(\alpha)$ where $\alpha$ is algebraic over $\mathbb{Q}$ with minimum polynomial
$a^{8}+6a^{4}+1=0.$
The automorphism is described in terms of its lift $f_{OY}$ to $X_{OY}$. The pullback of the coordinate $t$ is given by
$f_{OY}^{*}(t)=\frac{x+(\frac{1}{2}\cdot\alpha^{7}+3\cdot\alpha^{3}-\frac{1}{2}\cdot\alpha)\cdot y\cdot t+(\frac{1}{2}\cdot\alpha^{6}+\frac{5}{2}\cdot\alpha^{2})\cdot t^{6}+(-\frac{1}{2}\cdot\alpha^{6}-\frac{5}{2}\cdot\alpha^{2})\cdot t^{2}}{x\cdot t^{3}+(\frac{1}{2}\cdot\alpha^{7}+3\cdot\alpha^{3}-\frac{1}{2}\cdot\alpha)\cdot y-t^{5}+t}$
The equations of the pullback of $x$ and $y$ are far too large to display here. They can be computed from the file reproduction_final_result.jl.
The benefit of concrete equations is that we can now actually map points, reduce the surface to positive characteristic and specify a field of definition of surface and automorphism.
Corollary 1.2.
The surface $Y_{OY}$ is defined over the field of rational numbers $\mathbb{Q}$ and its automorphism of minimal entropy $\tilde{f}_{OY}$ is defined over the degree $8$ extension $\mathbb{Q}(\alpha)$.
Corollary 1.3.
For all prime numbers $p>2$ there exists an Enriques surface $Y_{p}$ over $\mathbb{F}_{p}$ and an automorphism $f_{p}\in\mathrm{Aut}(Y_{p}\times\mathrm{Spec}\,\mathbb{F}_{q})$ of entropy $\tau_{8}$ with $q=p^{n}$ and $n\leq 8$.
Proof.
Our model for $X_{OY}$ is defined over $\mathbb{Z}$ and has good reduction away from $2$. The Enriques involution is also defined over $\mathbb{Z}$. Since $p>2$, $\iota$ reduces to an Enriques involution $\iota_{p}$ on $X_{p}:=X_{OY}\times\mathrm{Spec}\,\mathbb{F}_{p}$. The quotient $Y_{p}:=X_{p}/\iota_{p}$ is defined over $\mathbb{F}_{p}$. Likewise $f_{OY}$ is defined over the localization $\mathcal{O}_{K}[2^{-1}]$ and reduces modulo primes of $\mathcal{O}_{K}$ over $p$. ∎
Question: Is $\tau_{8}$ is the minimum entropy of an Enriques surface in characteristic $p$?
1.1. Outline
 [^1] A first draft of this outline was created using ChatGPT4o. Then it was partially rewritten by the authors.
Section 2 briefly reviews K3 and Enriques surfaces, their Hodge structures, Torelli theorems, automorphism groups, and elliptic fibrations, as well as the role of lattice theory in classifying surface automorphisms. The next Section 3 focuses on identifying the Oguiso-Yu K3 surface $X_{OY}$ by searching a database for K3 surfaces with a matching transcendental lattice. The Enriques involution and automorphism group structure are analyzed, confirming that the correct surface has been found. Section 4 derives explicit equations by first representing $X_{OY}$ as a double cover $X_{h}\to\mathbb{P}^{1}\times\mathbb{P}^{1}$, then identifying an elliptic fibration and computing a Weierstrass model $X_{f_{1}}$. The automorphism of minimal entropy is then constructed in Section 5 explicitly, and its action on the Néron-Severi lattice is verified. In Section 6, we report on the framework for algebraic schemes used to carry out the computations and verify the end result. Among others, the computational treatment of sheaves, birational maps, specialization, intersection theory and pushforward of divisors is explained in the context of a scheme represented by an atlas of charts.
Acknowledgements
The first named author would like to thank Janko Böhm for sharing his insights. We would further like to thank Anne Frühbis-Krüger for collaboration and numerous helpful discussions on the schemes package in OSCAR, as well as Max Horn, Claus Fieker, and Tommy Hofmann for their enduring support in integrating our work into the OSCAR codebase. Finally, thanks go to Hans Schönemann for resolving various issues encountered with the Singular backend in OSCAR and again to Tommy Hofmann for providing numerous improvements on number-theoretic topics throughout the preparations of this work. This work was funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – Project-ID 286237555 – TRR 195.
2. Preliminaries
Since the universal cover of an Enriques surface is a K3 surface, we start with a brief review of K3 surfaces.
2.1. K3 surfaces
A complex K3 surface is a compact, connected, complex manifold $X$ with vanishing irregularity and trivial canonical bundle
$h^{1}(X,\mathcal{O}_{X})=0,\quad\omega_{X}\cong\mathcal{O}_{X}.$
The cup product gives the second singular cohomology group $H^{2}(X,\mathbb{Z})$ the structure of a quadratic $\mathbb{Z}$-lattice, which turns out to be even, unimodular, and of signature $(3,19)$. Such a form is unique up to isometry.
Since K3 surfaces are simply connected, linear, algebraic, and numerical equivalence agree and $\mathrm{Pic}(X)=\mathrm{NS}(X)=\mathrm{Num}(X)$. From the exponential sequence, we thus have an embedding $c_{1}\colon\mathrm{Pic}(X)\to H^{2}(X,\mathbb{Z})$ whose image is the subgroup of integral $(1,1)$-classes $H^{1,1}(X)\cap H^{2}(X,\mathbb{Z})$ by Lefschetz’ $(1,1)$-theorem. The smallest primitive sublattice of $H^{2}(X,\mathbb{Z})$ containing the period $H^{2,0}(X)$ is called the transcendental lattice $T(X)$. A K3 surface is not necessarily projective, but if it is, then $T(X)=\mathrm{NS}(X)^{\perp}$ in $H^{2}(X,\mathbb{Z})$.
Let $X$ be a complex K3 surface. The space of global $2$-forms $H^{0}(X,\omega_{X})\cong H^{2,0}(X)$ is $1$-dimensional and generated by a nowhere vanishing symplectic form $\sigma_{X}$. We denote by $\mathrm{Aut}_{s}(X)$ the kernel of the natural homomorphism
$\mathrm{Aut}(X)\to\mathrm{GL}(\mathbb{C}\sigma_{X}).$
Its elements are called symplectic automorphisms because they preserve the symplectic form $\sigma_{X}$. The other ones are called non-symplectic. For $G\leq\mathrm{Aut}(X)$ set $G_{s}=\mathrm{Aut}_{s}(X)\cap G$. Note that if $X$ is projective, then $\mathrm{Aut}(X)/\mathrm{Aut}_{s}(X)$ is a finite cyclic group.
Complex K3 surfaces are governed to a large extent by their Hodge structure. Let $X$ and $Y$ be two K3 surfaces. A Hodge-isometry is a linear map
$\phi\colon H^{2}(Y,\mathbb{Z})\to H^{2}(X,\mathbb{Z})$
which preserves the intersection form and the Hodge structures, i.e. $\phi_{\mathbb{C}}(H^{2,0}(Y))=H^{0,2}(X)$. The weak Torelli theorem states that $X$ and $Y$ are isometric, if and only if $H^{2}(X,\mathbb{Z})$ and $H^{2}(Y,\mathbb{Z})$ are Hodge-isometric.
To formulate the strong Torelli theorem, we need one more definition. The Hode-isometry $\phi$ is called effective, if it satisfies one of the following three equivalent conditions:
  1. (1)
    $\phi$ maps effective classes on $Y$ to effective classes on $X$,
  2. (2)
    $\phi$ maps Kähler classes to Kähler classes,
  3. (3)
    $\phi$ maps a single Kähler class to a Kähler class.
If $X$ is projective, then in (2) and (3) we may replace Kähler by ample.
Theorem 2.1 (Strong Torelli theorem).
Let $X,Y$ be K3 surfaces and
$\phi\colon H^{2}(Y,\mathbb{Z})\to H^{2}(X,\mathbb{Z})$
an effective Hodge isometry. Then there exists a unique isomorphism $f\colon X\to Y$ such that $\phi=f^{*}$.
Since the K3 surfaces covering Enriques surfaces are projective, we shall from now on assume that $X$ is projective and continue with the ample cone in place of the Kähler cone. The connected component of
$\{x\in\mathrm{NS}(X)\otimes\mathbb{R}\mid x^{2}>0\}$
which contains an ample class, is called the positive cone $\mathcal{P}_{X}$. Let
$\Delta(X)=\{x\in\mathrm{NS}(X)\mid x^{2}=0\}.$
By the Riemann-Roch theorem we have that for $r\in\Delta(X)$ either $r$ or $-r$ is effective. Set $\Delta^{+}(X)=\{r\in\Delta(X)\mid r\mbox{ is effective}\}$. Then $h\in\mathcal{P}_{X}$ is ample if and only if $h.r>0$ for all $r\in\Delta^{+}(X)$. More generally, the connected components of $\mathcal{P}_{X}\setminus\bigcup\{r^{\perp}\otimes\mathbb{R}\mid r\in\Delta(X)\}$ are called Weyl-chambers. The ample cone is one such chamber. The Weyl group $W(X)\subseteq O(\mathrm{NS}(X))$ is generated by the reflections into the hyperplanes $r^{\perp}$. It acts simply-transitively on the set of Weyl-chambers.
Since $\Delta^{+}(X)$ is often infinite, it requires some work to control the ample cone. Useful tools are Vinberg’s algorithm [Vinberg72] and the algorithms described in [Shimada24].
2.2. Enriques surfaces
A complex Enriques surface $Y$ is a compact complex manifold with $h^{1}(Y,\mathcal{O}_{Y})=0$ and $b_{2}(Y)=10$. Its canonical bundle $\omega_{Y}\neq\mathcal{O}_{Y}$, but $\omega_{Y}^{\otimes 2}\cong\mathcal{O}_{Y}$. Since the canonical bundle is $2$-torsion, it induces a $2:1$ cover $X\to Y$, which turns out to be the universal covering with $X$ a K3 surface. Since $h^{1,1}(Y)=b_{2}(Y)$, it follows from Lefschetz’ Theorem on $(1,1)$-classes and the Kodaira Embedding theorem that all Enriques surfaces are projective.
2.3. Elliptic Fibrations
An elliptic fibration on a smooth projective surface $S$ is a morphism $S\to C$ to a curve $C$ whose general fiber is an elliptic curve. If the map admits a section, we call the fibration Jacobian.
An elliptic fibration is called relatively minimal, if its fibers do not contain any $(-1)$-curves. The singular fibers of a relatively minimal elliptic fibration have been classified by Kodaira into various types and they can be computed from a Weierstrass model by Tate’s algorithm (see e.g. [Silverman94]). The dual graph of a reducible fiber is an extended ADE-diagram.
For K3 and Enriques surfaces it is known that $C=\mathbb{P}^{1}$. Elliptic fibrations on Enriques surfaces are never Jacobian; on a K3 surface they may or may not be Jacobian. If it is, then its generic fiber admits a Weierstrass model
$y^{2}+a_{1}xy+a_{3}y=x^{3}+a_{4}x^{2}+a_{2}x+a_{6}$
with $a_{i}=a_{i}(t)\in\mathbb{C}(t)$ of degree at most $2i$ (and further conditions assuring that it is smooth and minimal).
2.4. Lattices
Let $L$ be a quadratic integer lattice and $f\in O(L)$ an isometry. The invariant and co-invariant lattices of $f$ are
$L^{i}=\{x\in L\mid f(x)=x\}\qquad L_{i}:=(L^{i})^{\perp}.$
We follow the geometric convention that ADE-root lattices are negative definite.
3. The Hodge- and lattice theory of the surface
In this section we deduce as much information as we can from the Hodge-theoretic model of the sought for Enriques surface $Y_{OY}$ and its automorphism of minimal entropy, in order to guide our search for explicit equations.
3.1. The construction of Oguiso and Yu
We review the construction of Oguiso and Yu in [OguisoYu20, Theorem 8.1 and proof]. Denote by $Y_{OY}$ the Oguiso-Yu Enriques surface, by $\widetilde{f}_{OY}$ an automorphism of entropy $h(\widetilde{f}_{OY})=\log\tau_{8}$, by $\iota_{OY}$ the covering involution of the universal cover $X_{OY}\to Y_{OY}$ of $Y_{OY}$ by the K3 surface $X_{OY}$, and by $f_{OY}\colon X_{OY}\to X_{OY}$ a lift of $\widetilde{f}_{OY}$. Oguiso and Yu construct
  1. a)
    an even unimodular lattice $\Lambda_{K3}$ of signature $(3,19)$
  2. b)
    an isometry $i^{\prime}\in O(\Lambda_{K3})$ with $i^{\prime 2}=1$ and invariant lattice $\Lambda_{K3}^{i^{\prime}}\cong U(2)\oplus E_{8}(2)$,
  3. c)
    an isometry $f^{\prime}\in O(\Lambda_{K3})$ with $i^{\prime}\circ f^{\prime}=f^{\prime}\circ i^{\prime}$ and spectral radius $\lambda(f^{\prime})=\tau_{8}$.
Next, it is proven that there exists a Hodge structure of weight two on $\Lambda_{K3}$, preserved by both $i^{\prime}$ and $f^{\prime}$. Moreover, these two morphisms preserve the same Weyl-chamber and in particular, they are effective. Then it follows from the surjectivity of the period map and the strong Torelli theorem that there exist
  1. 1)
    a K3 surface $X_{OY}$,
  2. 2)
    an Enriques involution $\iota_{OY}\in\mathrm{Aut}(X_{OY})$,
  3. 3)
    an automorphism $f_{OY}\in\mathrm{Aut}(Y_{OY})$,
  4. 4)
    a marking $\eta\colon H^{2}(X_{OY},\mathbb{Z})\xrightarrow{\sim}\Lambda_{K3}$ such that $\eta\circ\iota_{OY}^{*}=i^{\prime}\circ\eta$ and $\eta\circ f_{OY}^{*}=f^{\prime}\circ\eta$.
Finally, $f_{OY}$ descends to $Y_{OY}=X_{OY}/\iota_{OY}$ since it commutes with the Enriques involution $\iota_{OY}$.
All we get explicitly from this construction are a), b), and c) in the form of three $22\times 2$ 22 integer matrices representing the gram matrix of $\Lambda_{K3}\cong\mathbb{Z}^{22}$ and the isometries $i^{\prime}$ and $f^{\prime}$. It is our task to find 1) to 4).
3.2. The transcendental lattice
By [OguisoYu20, Theorem 8.1] the transcendental lattice $T(X_{OY})$ is isometric to $L(4)$ with $L$ an odd unimodular lattice of signature $(2,2)$. Furthermore, $f_{OY}$ acts with order $8$ on $T(X_{OY})$. Since $\varphi(8)=4$ (the Euler-phi-function), the characteristic polynomial of $f_{OY}|T(X_{OY})$ is $x^{4}+1$. The complexification $T(X_{OY})\otimes\mathbb{C}$ contains the period
$H^{2,0}(X_{OY})\cong H^{0}(X_{OY},\Omega^{2}_{X_{OY}})=\mathbb{C}\sigma_{OY},$
which is preserved by $f_{OY}$. Hence, it must be one of the four eigenspaces of $f_{OY}|T(X_{OY})\otimes\mathbb{C}$. Moreover, we know that $\sigma_{OY}.\overline{\sigma_{OY}}>0$. This leaves us the possibility of only two complex conjugate eigenspaces which could contain $\sigma_{OY}$. Which one it is, is decided by the eigenvalue. We see that the action of $f_{OY}$ on $H^{2}(X_{OY},\mathbb{Z})$ and the eigenvalue of the eigenspace $H^{2,0}(X)\subseteq T(X_{OY})\otimes\mathbb{C}$ determine the Hodge structure of $X_{OY}$. The next theorem shows that under suitable assumptions the isomorphism class of the action on $T(X)$ and the eigenvalue determine the Hodge structure up to isomorphism.
Theorem 3.1.
[brandhorst:purely-nonsymplectic, Theorem 1.2] Let $X_{i}$, $i=1,2$ be two complex K3 surfaces and $f_{i}\in\mathrm{Aut}(X_{i})$ automorphisms with $f_{i}^{*}\sigma_{i}=\zeta_{n}\sigma_{i}$ where $\zeta_{n}=\exp(2\pi i/n)$, $n\in\mathbb{N}$ and $\sigma_{i}\in H^{0}(X_{i},\Omega_{X_{i}}^{2})$ is a non-zero holomorphic $2$-form on $X_{i}$. Let $I_{i}$ be the kernel of the natural map given by
$\mathbb{Z}[\zeta_{n}]\to O(\mathrm{NS}(X_{i})^{\vee}/\mathrm{NS}(X_{i})),\quad\zeta_{n}\mapsto f_{i}^{*}.$
If $f_{1}$ is of finite order, then $X_{1}\cong X_{2}$ if and only if the ideals $I_{1}$ and $I_{2}$ are equal.
At this point it is plausible to search for some K3 surface with a non-symplectic automorphism of order $8$, having the same transcendental lattice as $X_{OY}$.
A very general K3 surface with a purely non-symplectic automorphism of order $8$ has $2$-elementary transcendental lattice and $\mathrm{rank}\,T(X)=4r$, a multiple of $4$. Thus we have to look for a special member in a family of K3 surfaces with a non-symplectic automorphism of order $8$. One way to force extra algebraic cycles (hence forcing $r=1$) is to require the existence of additional symplectic automorphisms. We can therefore hope to find $X_{OY}$ among the set of K3 surfaces $X$ admitting a group $G\leq\mathrm{Aut}(X)$ with $G/G_{s}\cong C_{8}$ and its subgroup of symplectic automorphisms $G_{s}\neq 1$.
The database [BrandhorstHofmann23] consists of a complete set of representatives of pairs $(X,G)$ where $X$ is a K3 surface and $G\leq\mathrm{Aut}(X)$ up to deformation and isomorphism. Each entry of the database corresponds to a moduli space of such pairs $(X,G)$. The moduli space is encoded by an even unimodular lattice $\Lambda_{K3}$ and a (satured and effective) group $G^{\prime}\leq O(\Lambda_{K3})$, given in terms of matrices. Via a marking (which we do not have explicit access to) $\Lambda_{K3}$ can be identified (non-uniquely!) with $H^{2}(X,\mathbb{Z})$ and $G^{\prime}$ with the image of $G\to O(H^{2}(X,\mathbb{Z}))$. For a generic member $X$ of the moduli space this allows to compute the isomorphism classes of $\mathrm{NS}(X)$, $T(X)$ and various further sublattices associated to the action of $G$ and $G_{s}$.
The database contains $108$ pairs $(\Lambda_{K3},G^{\prime})$ with $G_{s}\neq 1$ and $|G/G_{s}|=8$. Among these, $66$ have $\mathrm{rank}\,T(X)=4$. They have $\det T(X)\in\{4,16,64,256,324\}$. We are interested in $\det T(X)=256$. There are $18$ such database entries and all of them have $T(X)\cong T(X_{OY})$. They satisfy that $|G_{s}|\in\{4,8,16,32\}$. We will take $G_{s}$ as large as possible in the expectation that additional symmetry helps us to find the equations for $X$ and algebraic cycles in $\mathrm{NS}(X)$ and in the hope that $G$ contains the Enriques involution $\iota_{OY}$ we seek.
Proposition 3.2.
Let $X$ be a complex K3 surface and $G\leq\mathrm{Aut}(X)$ such that $|G|=256$ and $G/G_{s}\cong C_{8}$. Then the pair $(X,G)$ is unique up to isomorphism and given by the entry 40.8.0.1 in the database [BrandhorstHofmann23]. Moreover, $G_{s}\cong(Q_{8}*Q_{8})$ and $X\cong X_{OY}$.
Proof.
Only the entry numbered 40.8.0.1 has a group action by a group $G$ of order $256$ and $G/G_{s}\cong C_{8}$. The corresponding moduli space consists of a single point and therefore the pair $(X,G)$ with the given conditions is unique up to isomorphism. Using the Hodge-theoretic model in the database, one computes that $T(X)$ is in the same genus as $T(X_{OY})$. The rescaled lattices $T(X)(1/4)$ and $T(X_{OY})(1/4)$ are both odd unimodular lattices of signature $(2,2)$. By the classification of indefinite odd unimodular lattices in terms of their signature (see e.g. [Serre78, Chapter V §2.2 Theorem 4]), they are abstractly isomorphic.
Let $gG_{s}$ be a generator of $G/G_{s}$. We want to apply Theorem 3.1 with $X_{1}=X_{h}$, $X_{2}=X_{OY}$, $f_{1}=g^{k_{1}}$ and $f_{2}=f_{OY}^{k_{2}}$ for $k_{1}$ and $k_{2}$ suitable odd integers such that $f_{i}$ acts by $\zeta_{8}$ on the space of holomorphic $2$-forms $H^{2,0}(X_{i})$ for $i=1,2$. Since $T(X)\cong T(X_{OY})$ as lattices, the corresponding ideals $I_{i}\leq\mathbb{Z}[\zeta_{8}]$ are both of norm $N(I_{i})=\det T(X)=4^{4}$. The prime $2$ is completely ramified in the extension $\mathbb{Q}(\zeta_{8})/\mathbb{Q}$. Therefore, there is a unique ideal $I\leq\mathbb{Z}[\zeta_{8}]$ of norm $4^{4}=\det T(X)$. We see that $I_{1}=I_{2}$ and by Theorem 3.1 $X\cong X_{OY}$. ∎
Next, we need to search for clues where to find the Enriques involution $\iota_{OY}$ giving $Y_{OY}$. By chance, it turns out that $\iota_{OY}$ is contained (up to conjugacy) in $G$.
Remark 3.3.
Let $X$ be a K3 surface and $i,j\in\mathrm{Aut}(X)$ be two Enriques involutions. The Enriques surfaces $X/i$ and $X/j$ are isomorphic if and only if $i$ and $j$ are $\mathrm{Aut}(X)$-conjugate [ohashi]. A necessary condition for this is that the algebraic coinvariant lattices $\mathrm{NS}(X)_{i}$ and $\mathrm{NS}(X)_{j}$ are isomorphic. Note that $\mathrm{NS}(X)_{i}$ of $i$ does not contain any vectors $x$ with $x^{2}=-2$, since $i$ preserves the nef cone.
By a direct computation with the Hodge-theoretic model provided by Oguiso and Yu, one finds that $\mathrm{NS}(X_{OY})_{\iota_{OY}}$ is of determinant $2^{10}$ and lies in the genus $\mathrm{II}_{0,8}(2^{-2}4^{-4}_{4})$. An application of Kneser’s neighbor method [Kneser02, Kapitel IX] (see [Voight23] for a survey), as shipped with Oscar [oscar], shows that this genus contains a total of $3$ isometry classes, two of which have maximum $-2$ and one maximum $-4$. The unique one of maximum $-4$ must be the one isometric to $\mathrm{NS}(X_{OY})_{\iota_{OY}}$.
Lemma 3.4.
Let $(X,G)$ be the symmetric K3 surface of id 40.8.0.1. Then $G$ contains up to conjugacy exactly two Enriques involutions $j_{1}$ and $j_{2}$. Their coinvariant lattices lie in the genera
$\mathrm{NS}(X)_{j_{1}}\in\mathrm{II}_{0,8}(2^{-2}4^{-4}_{4})\mbox{ and }\mathrm{NS}(X)_{j_{2}}\in\mathrm{II}_{0,8}(2^{-4}4^{-4}_{4}).$
The involution $j_{1}$ is contained in the center $C(G)\cong C_{2}\times C_{2}$, but $j_{2}$ is not.
Proof.
An involution $j\in G$ is an Enriques involution if and only if its fixed lattice $F:=H^{2}(X,\mathbb{Z})^{j}$ is isometric to $E_{10}(2)$ if and only if $F(1/2)$ is an even unimodular lattice (of signature $(1,9)$) [Keum90]. Since the database contains abstractly the action of $G$ on $H^{2}(X,\mathbb{Z})$, the conditions can be checked on a complete set of representatives of the conjugacy classes of involutions in $G$, yielding that $G$ contains two Enriques involutions, and that one of them is in the center. The coinvariant lattices and their genus can be computed explicitly. These computations are carried out in the file lattice_computations.jl. ∎
Proposition 3.5.
Let $X:=X_{OY}$ and $j\in\mathrm{Aut}(X)$ an Enriques involution with
$\mathrm{NS}(X)_{j}\in\mathrm{II}_{0,8}(2^{-2}4^{-4}_{4}).$
Then $j$ is $\mathrm{Aut}(X)$-conjugate to $\iota_{OY}$.
Proof.
Using the theory of Nikulin and Miranda–Morrison [nikulin, miranda_morrison:embeddings] and its OSCAR implementation [BrandhorstHofmann23, brandhorst-veniani] one checks that any two primitive sublattices $F,F^{\prime}\leq\mathrm{NS}(X)$ with $F\cong F^{\prime}\cong\mathrm{NS}(X)_{j}$ are in the same $O(\mathrm{NS}(X))$-orbit.
Let $O(\mathrm{NS}(X),j)$ be the centraliser of $j$ in $O(\mathrm{NS}(X))$. A direct (but non-trivial) computation based on Muller’s Oscar package [muller:quadform_and_isom, BrandhorstHofmann23] shows that the composition
$O(\mathrm{NS}(X),j)\to O(\mathrm{NS}(X))\to O(\mathrm{NS}(X)^{\vee}/\mathrm{NS}(X))$
is surjective. It follows from Shimada-Veniani [shimada-veniani, Theorem 3.1.9] that any Enriques involution $k\in\mathrm{Aut}(X)$ with $\mathrm{NS}(X)_{k}\cong\mathrm{NS}(X)_{j}$ is $\mathrm{Aut}(X)$-conjugate to $j$. Take $k=\iota_{OY}$ to obtain the statement. See the file lattice_computations.jl for the computations. ∎
Proposition 3.6.
Let $(X,G)$ be the symmetric K3 surface of id 40.8.0.1 and denote by $j$ the unique Enriques involution in the center of $G$. Then $X/j\cong Y_{OY}=X_{OY}/\iota_{OY}$.
Proof.
By Lemma 3.4 $\mathrm{NS}(X_{OY})_{\iota_{OY}}\cong\mathrm{NS}(X)_{j}$. The statement of this proposition follows with Proposition 3.5 and Remark 3.3. ∎
To aid the search for a projective model of $X$ and $G$, we look at $G$-invariant ample divisors on $X$ and their properties. Note that since $G$ is finite, there exists a $G$-invariant ample class. Since $\mathrm{rank}\,\mathrm{NS}(X)^{G}=1$, the invariant lattice is spanned by a unique primitive ample class $h\in\mathrm{NS}(X)^{G}$. We have $h^{2}=4$. Furthermore, we know that $h$ is basepoint free, because, by a direct calculation, the set
$\{f\in\mathrm{NS}(X)\mid f^{2}=0,h.f=1\}=\emptyset$
is empty and hence $h$ induces a morphism $\phi:X\to\mathbb{P}^{3}$, cf. [SaintDonat74]. In general, the ample class $h$ can belong to one of the following three types according to [SaintDonat74]:
  1. A)
    very ample: $\phi$ is an embedding
  2. H)
    hyperelliptic: $\phi$ is a $2:1$ cover onto a quadric hypersurface $Q\cong\mathbb{P}^{1}\times\mathbb{P}^{1}$ branched over a divisor of bidegree $(4,4)$.
  3. U)
    unigonal: the generic fiber of $\phi$ is an elliptic curve and its image is a twisted cubic.
Proposition 3.7.
Let $(X,G)$ be the symmetric K3 surface of id 40.8.0.1. The (unique) $G$-invariant ample class $h\in\mathrm{NS}(X)^{G}$ is basepoint free hyperelliptic.
Proof.
We have seen already that $h$ is basepoint free. Now one computes that
$\{e\in\mathrm{NS}(X)\mid e^{2}=0,h.e=2\}=\{e_{1},e_{2}\}.$
[…]
Remark 3.8.
The ample class $h\in\mathrm{NS}(X)^{G}$ is invariant under the Enriques involution $j\in G$. Hence $h$ descends to an ample class $h_{Y}$ of degree $2$ on $Y_{OY}\cong X/j$. The elliptic divisors $e_{1},e_{2}\in\mathrm{NS}(X)$ with $e_{1}+e_{2}=h$ induce elliptic fibrations on $X$. Since they are invariant under $j$, they descend to $Y$ and induce elliptic fibrations on $Y$. The morphism corresponding to the sum of the two half fibers $\pi_{*}(e_{i})$ gives the so called Horikawa model of non-special Enriques surfaces. Their geometry is described for instance in [BarthHulekPetersVen04, VIII Proposition 18.1].
Let us compute the type of the elliptic fibrations induced by $e_{1}$ and $e_{2}$.
Proposition 3.9.
Let $i\in\{1,2\}$. The elliptic fibration induced by $e_{i}$ on $X$ has $8$ reducible fibers of type $A_{1}$ and Mordell-Weil group $\mathbb{Z}/2\mathbb{Z}\oplus\mathbb{Z}^{8}$. The stabiliser of $e_{i}$ in $G$ is of order $128$ and its orbit is $\{e_{1},e_{2}\}$. The elliptic fibration induced by $e_{i}$ on $Y=X/j$ has $4$ simple $A_{1}$ fibers.
Proof.
The description of the stabilisers and the orbit follows from the proof of Proposition 3.7 as follows: Since $h$ is invariant under $G$, so is the set
$\{f\in\mathrm{NS}(X)\mid f^{2}=0,h.f=2\}=\{e_{1},e_{2}\}.$
Since the invariant lattice is of rank $1$, $G$ must exchange $e_{1}$ and $e_{2}$ and the stabiliser of $e_{i}$ is of order $G/2=128$. Recall from [shioda] that the trivial lattice $\mathrm{Triv}(f)$ of an elliptic fibration with fiber $f\in\mathrm{NS}(X)$ is the sublattice of $\mathrm{NS}(X)$ spanned by the zero section and fiber components. The Mordell-Weil group is isomorphic to $\mathrm{NS}(X)/\mathrm{Triv}(e_{i})$. To compute the trivial lattice, one chooses any $s\in\mathrm{NS}(X)$ with $s^{2}=-2$ and $s.f=1$. Let $R$ be the root sublattice of $\langle f,s\rangle^{\perp}\subseteq\mathrm{NS}(X)$. Then the trivial lattice is isomorphic to $\mathbb{Z}\langle s,f\rangle\oplus R$. The irreducible root sublattices of $R$ give the ADE types of the singular fibers. For how to compute the fiber types on the Enriques surface $Y$ from the class of $e_{i}$ see [brandhorst-gonzalez, Prop. 5.8]. ∎
Remark 3.10.
With the methods described in [brandhorst-gonzalez, Section 5] and the Hodge and lattice theoretic model of $(X,G)$, one computes the following: Let $Y_{OY}$ be the Oguiso-Yu Enriques surface. Modulo $\mathrm{Aut}(Y_{OY})$ it contains exactly $12$ elliptic fibrations. The following table contains the ADE types of the reducible fibers and whether they are double, the count of $\mathrm{Aut}(Y_{OY})$ orbits with the same fiber types, and the ramification index of the fibration in the forgetful map $\mathcal{P}^{0}\colon\mathcal{M}_{En,0}\to\mathcal{M}_{En}$ of the moduli space of elliptic Enriques surface to the moduli space of Enriques surfaces.
simple fibers double fibers count ramification index
$1A_{1}+A_{3}$ 1 128
$2A_{1}+A_{3}$ 1 128
$3A_{1}+A_{3}$ 2 64
$6A_{1}$ 1 48
$4A_{1}$ 1 48
$4A_{1}$ $2A_{1}$ 1 24
$2A_{1}+D_{4}$ 2 8
$2A_{1}+D_{4}$ $A_{1}$ 1 4
$2A_{1}$ $A_{3}+A_{1}$ 1 2
$4A_{1}$ $A_{3}$ 1 1
As a sanity check we calculate that the weighted count of isomorphism classes of elliptic fibrations equals
$128+128+2\cdot 64+48+48+24+2\cdot 8+4+2+1=527,$
which agrees with the degree of the forgetful map $\mathcal{P}^{0}$ – as it should.
Besides a projective model for $X_{OY}$ and an elliptic fibration, we also seek explicit generators for $\mathrm{NS}(X_{OY})$. A common approach is to search for smooth rational curves of small degree. The lattice side will tell us up to which degree we have to search:
Proposition 3.11.
Let $(X,G)$ be the symmetric K3 surface of id 40.8.0.1 and $h$ be the unique $G$-invariant ample class of $X$. Denote by $D_{i}$ the set of smooth rational curves $c$ on $X$ of degree $h.c=i$. Then $|D_{1}|=32$, $\mathrm{rank}\,\langle D_{1}\rangle=14$ and $D_{1}$ forms a single $G$-orbit. The set $D_{2}$ is of cardinality $160$ and splits into three $G$-orbits of sizes $32$, $64$, and $64$. Altogether, $D_{1}$ and the orbit of size $32$ in $D_{2}$ span a sublattice $L$ of $\mathrm{NS}(X)$ of rank $18$, determinant $2^{2}4^{4}$ and index $[\mathrm{NS}(X):L]=2$.
Proof.
The classes in $\mathrm{NS}(X)$ of the smooth rational curves of $X$ are the facets of the nef cone, which itself is a fundamental chamber for the action of the Weyl group generated by the $(-2)$-hyperplanes. We use Vinberg’s algorithm[Vinberg72] to compute the facets of smallest distance to $h$ in hyperbolic space:
$D_{1}=\{x\in\mathrm{NS}(X)\mid h.x=1,x^{2}=-2\},$
$D_{2}=\{y\in\mathrm{NS}(X)\mid h.y=2,y^{2}=-2,\forall x\in D_{1}:x.y\geq 0\}.$
These sets can be computed by solving a suitable close vector problem. See the file lattice_computations.jl for the computations. ∎
At this point we are given on the one hand $(\Lambda_{K3},\iota^{\prime},f^{\prime})$ corresponding to $X_{OY}$, $\iota_{OY}$ and $f_{OY}$. On the other hand from the database entry 40.8.0.1 we get $L_{K3},G^{\prime},j^{\prime},h^{\prime}$ corresponding to $X,G$, and $h$. Both sides come with implicitly defined Hodge-structures and Weyl-chambers. Since $X_{OY}\cong X$, we need to reconcile these two points of views, i.e. we need to find an explicit effective Hodge isometry $\Lambda_{K3}\to L_{K3}$ that conjugates $\iota^{\prime}$ to $j^{\prime}$. Ideally, it should map $f^{\prime}$ to a matrix with small coefficients in our choice of basis for $\mathrm{NS}(X_{OY})\otimes\mathbb{Q}$ from $D_{1}\cup D_{2}$ (or rather $f^{\prime}(h).h$ should be small).
Given the existing computational infrastructure it is easier to compute generators for the image of $\mathrm{Aut}(Y)\to\mathrm{Num}(Y)$ (where $Y=X/j$ and $\mathrm{Num}(Y)(2)$ is identified with $L_{K3}^{j}$) using Borcherds’ method as described in [brandhorst_shimada:tautaubar] and search for an automorphism with entropy $\log\tau_{8}$. This has the additional merit of possibly factoring $f_{OY}$ into isometries of zero entropy and hence a simpler geometry.
In the following proposition we collect our findings:
Proposition 3.12.
Let $Y=X/j$ be the Oguiso-Yu Enriques surface where $(X,G)$ is the entry 40.8.0.1 from [BrandhorstHofmann23]. Then $X$ carries an automorphism $f$ of entropy $\log\tau_{8}$ and commuting with $j$. It factors as $f=l\circ g$ where $g\in G$ and $l$ preserves an elliptic fibration $\phi_{l}\colon X\to\mathbb{P}^{1}$ with two $\widetilde{A}_{1}$ fibers and two $\widetilde{A}_{3}$ fibers. The Mordell-Weil group of $\phi_{l}$ is $\mathbb{Z}^{8}$ and this fibration descends to $Y$ with a reducible fiber of type $A_{1}$ and a reducible fiber of type $A_{3}$.
We have collected enough information to guide the search for equations of the symmetric K3 surface $(X,G)$, for the elliptic fibration on $X$ given by the isotropic class $e_{1}$ and for generators of its Néron-Severi lattice $\mathrm{NS}(X)$. Once we have these, it remains to connect them by an effective Hodge isometry to the abstract K3 surface found in the database.
4. Construction of the Enriques surface
4.1. The first model
[…]
(1) $\Phi\colon X_{h}\overset{2:1}{\twoheadrightarrow}Q\cong\mathbb{P}^{1}\times\mathbb{P}^{1}\hookrightarrow\mathbb{P}^{3}$
[…]
Eventually, we find that there is a projective representation $\rho\colon G\to\mathbb{P}GL(3;\mathbb{Q}(\zeta_{8}))$ over the cyclotomic field generated by a primitive $8$-th root of unity $\zeta_{8}$. The representation $\rho$ is not faithful and its kernel is generated by an involution $\eta\in G$. The image $\tilde{G}\subset\mathbb{P}GL(3;\mathbb{Q}(\zeta_{8}))$ of this representation is the projective matrix group generated by [^2] Note that in accordance with the OSCAR convention matrices act from the right throughout this paper and a “vector” is a row-vector unless otherwise specified.
(2) $\widetilde{g}_{1}=\begin{pmatrix}0&1&0&0\\ -1&0&0&0\\ 0&0&1&0\\ 0&0&0&-1\end{pmatrix},\quad\widetilde{g}_{2}=\begin{pmatrix}0&0&1&0\\ 0&0&0&1\\ \zeta_{8}&0&0&0\\ 0&-\zeta_{8}^{3}&0&0\end{pmatrix}.$
The center of $\widetilde{G}$ is of order $2$ and generated by $\widetilde{i}:=\widetilde{g}_{1}^{2}$.
The relative invariants of $\rho$ are given by two polynomials
$q(x,y,z,w)=xy+zw,\quad b(x,y,z,w)=x^{4}+y^{4}-iz^{4}-iw^{4}.$
The expectation now is that $q$ defines the quadric $Q$ from the system (1), embedded as $\mathbb{P}^{1}\times\mathbb{P}^{1}\hookrightarrow\mathbb{P}^{3}$ via
$[(s:t),(u:v)]\mapsto(x:y:z:w)=(su:tv:tu:-sv)$
and that the vanishing locus of $b$ is the branch locus $B$ of $\Phi$. We thus construct a $2:1$ covering from $b$ by setting
(3) $\lambda^{2}=b(1,y,z,w)$
for the affine coordinate $\lambda$ of the bundle $\mathbb{P}(\mathcal{O}\oplus\mathcal{O}(-2))$ over $Q$. Extending this in the obvious way to all other affine charts, we obtain a surface
${X_{h}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathbb{P}(\mathcal{O}\oplus\mathcal{O}(-2))\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${Q\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\pi_{1}}$$\scriptstyle{\pi_{2}}$${\mathbb{P}^{1}}$${\mathbb{P}^{1}}$
with two natural projections to $\mathbb{P}^{1}$ induced from the rulings of $Q\cong\mathbb{P}^{1}\times\mathbb{P}^{1}$. This construction is carried out in minimum_entropy_surface.jl, lines 1–72. By direct calculation we verify that the curve $B$ is smooth and, hence, by [BarthHulekPetersVen04, Chapter V Section 23], $X_{h}$ is a K3 surface.
Let $\widetilde{j}\in C(\widetilde{G})$ be the involution in the center of $\widetilde{G}$. It is given by $\widetilde{j}(x,y,z,w)=(-x,-y,z,w)$. It has $4$ isolated fixed points and none of them lies on the branch curve $B$. By [BarthHulekPetersVen04, Chapter V Section 23] the involution $\widetilde{j}$ lifts to a fixed-point-free involution $j\in G$ of $X_{h}$ and $Y:=X_{h}/j$ is the Horikawa model of a so called non-special Enriques surface. Further, the two rulings of $Q$ define elliptic fibrations $X_{h}\to Q\to\mathbb{P}^{1}$.
It is easy to see that both $\pi_{1}$ and $\pi_{2}$ induce elliptic-fibrations on $X_{h}$: The fiber $\pi_{1}^{-1}(\{p\})=\{p\}\times\mathbb{P}^{1}$ of any point $p\in\mathbb{P}^{1}$ is a projective line. The preimage of that space in $X_{h}$ is a curve $C$ with a generically $2:1$-projection $C\to\mathbb{P}^{1}$, ramified over the points where $b=0$. For a generic choice of $p$ the curve $C$ is smooth and there are four simple ramification points. Then the Hurwitz formula for morphisms of smooth curves asserts that in fact $g(C)=1$. In the following we will denote the composed maps $X_{h}\to\mathbb{P}^{1}$ by $\pi_{1}$ and $\pi_{2}$, as well.
Following Proposition 3.2 we know that we have found the pair $(X,G)$ of id 40.8.0.1 as soon as we show that $G/G_{s}\cong C_{8}$, i.e. that $G$ acts with order $8$ on $H^{2,0}(X_{h},\mathbb{C})=H^{0}(X_{h},\Omega_{X_{h}}^{2})$. We quickly verify this via explicit calculation as follows.
Lemma 4.1.
A global non-vanishing holomorphic $2$-form $\sigma_{X_{h}}\in H^{0}(X_{h},\Omega^{2}_{X_{h}})$ is given by the natural extension of
$\frac{1}{2\lambda}\operatorname{d}\!t\wedge\operatorname{d}\!v=\frac{1}{4(v^{4}-i)t^{3}}\operatorname{d}\!\lambda\wedge\operatorname{d}\!v=\frac{1}{4(t^{4}-i)v^{3}}\operatorname{d}\!t\wedge\operatorname{d}\!\lambda$
in the defining chart of (3) and the affine coordinates $(t,v)$ of $\mathbb{P}^{1}\times\mathbb{P}^{1}$.
Proof.
This is a slightly tedious, but straightforward computation: The automorphism on $\mathbb{P}^{1}\times\mathbb{P}^{1}$ induced by $\tilde{g}_{2}$ is given by
$\tilde{g}_{2}\colon(t,v)\mapsto\left(\frac{1}{\zeta_{8}t},\zeta_{8}^{3}v\right)$
in the standard affine coordinates $(t,v)$ of the chart $\{s\neq 0\}\cap\{u\neq 0\}$ of $\mathbb{P}^{1}\times\mathbb{P}^{1}$. The defining equation (3) pulls back to
(4) $\lambda^{2}=1+t^{4}v^{4}-it^{4}-iv^{4}$
in these coordinates and we can lift $\tilde{g}_{2}$ to an automorphism of $X_{h}$ by setting
(5) $\tilde{g}_{2}^{\prime}\colon(\lambda,t,v)\mapsto\left(\frac{\zeta_{8}\lambda}{t^{2}},\frac{1}{\zeta_{8}t},\zeta_{8}^{3}v\right).$
Indeed, for these choices we get
$\left(\tilde{g}_{2}^{\prime}\right)^{*}(1+t^{4}v^{4}-it^{4}-iv^{4})$ $=$ $\left(1+\frac{v^{4}}{t^{4}}+i\frac{1}{t^{4}}+iv^{4}\right)$
$=$ $i\frac{1}{t^{4}}\left(-it^{4}-iv^{4}+1+t^{4}v^{4}\right)$
$=$ $i\frac{1}{t^{4}}\lambda^{2}=\left(\frac{\zeta_{8}\lambda}{t^{2}}\right)^{2}=\left(\tilde{g}_{2}^{\prime}\right)^{*}(\lambda^{2}).$
But then
$\left(\tilde{g}_{2}^{\prime}\right)^{*}\sigma_{X_{h}}=\frac{t^{2}}{2\zeta_{8}\lambda}\cdot\left(\frac{-1}{\zeta_{8}t^{2}}\operatorname{d}\!t\right)\wedge\left(\zeta_{8}^{3}\operatorname{d}\!v\right)=-\zeta_{8}\cdot\sigma_{X_{h}}$
so that the lift of $\tilde{g}_{2}$ clearly acts with order $8$ on the subspace generated by $\sigma_{X_{h}}$.
Alternatively, one may argue as follows: By direct calculation $(\tilde{g}_{2}^{\prime})^{4}(\lambda,t,v)=(\lambda,t,-v)$. Hence it has a fixed curve given by $v=0$ and must be non-symplectic, because a finite order symplectic automorphism has only isolated fixed points. ∎
It then follows from Proposition 3.2 that $X_{OY}\cong X\cong X_{h}$ and from Proposition 3.6 that $X_{h}/j=:Y$ is isomorphic to the Oguiso-Yu Enriques surface $Y_{OY}$ of minimum entropy.
4.2. Search for sections
So far we have found the surface $X_{h}\cong X_{OY}$, the Enriques involution $\iota_{OY}$, and hence $Y_{OY}=X_{OY}/\iota_{OY}$. To hunt for equations of the automorphism $f_{OY}$ of minimum entropy, we need to find generators of the Néron-Severi lattice $\mathrm{NS}(X_{h})$ and a Weierstrass model $X_{e_{1}}$ of the elliptic fibration $\pi_{1}\colon X_{h}\to\mathbb{P}^{1}$; in particular a section.
Proposition 3.11 tells us that to generate a finite index sublattice of $\mathrm{NS}(X_{h})$, we have to search for smooth rational curves of degrees $1$ and $2$ contained in $X_{h}$. Since $h=e_{1}+e_{2}$ (with $e_{i}$ the class of a fiber of either one of the two elliptic-fibrations $\pi_{i}$) the smooth rational curves $C$ of degree $1=[C].h$ are contained in singular fibers of $\pi_{1}$ or $\pi_{2}$, hence easy to write down. Thus it remains to search for rational curves $D\subseteq X_{h}$ of degree $2=[D].h=[D].(e_{1}+e_{2})$. There are two such cases which can occur. Either $[D].e_{1}=0$ and $[D].e_{2}=2$ or vice versa and hence $D$ is contained in one of the fibers. Or –and this is the other case– $[D].e_{1}=[D].e_{2}=1$, i.e. $D$ is a section of both fibrations.
In order to find sections $D$ of the maps $\pi_{i}\colon X_{h}\to\mathbb{P}^{1}$ as divisors on $X_{h}$, we reason as follows. Write $e_{i}=\Phi^{*}\epsilon_{i}$ for some divisors $\epsilon_{i}$ on $Q$. Then $1=e_{i}.D=\Phi^{*}\epsilon_{i}.D=\epsilon_{i}.\Phi_{*}D$ for $i=1,2$ shows that $\Phi_{*}D\subset\mathbb{P}^{1}\times\mathbb{P}^{1}$ is the graph of an isomorphism $\mathbb{P}^{1}\to\mathbb{P}^{1}$. Further $\Phi^{*}\Phi_{*}D\neq D$, since otherwise
$e_{i}.D=(\Phi^{*}\epsilon_{i}).(\Phi^{*}\Phi_{*}D)=\deg\Phi\cdot(\epsilon_{i}.(\Phi_{*}D))=2\cdot(\epsilon_{i}.(\Phi_{*}D))$
would be even. Hence $\Phi^{*}\Phi_{*}D$ must split into two components $D_{1}+D_{2}$. This leads us to the following Ansatz. Let
$\Delta_{a}=\{((s:t),(s:a\cdot t)\in\mathbb{P}^{1}\times\mathbb{P}^{1}\mid(s:t)\in\mathbb{P}^{1}\},$
be a “tilted diagonal”. If we restrict the defining equation $b$ of the branch locus to $\Delta_{a}$, then it might split as a square, depending on $a$
$b(x,y,z,w)|_{\Delta_{a}}=b(s^{2},a\cdot t^{2},st,-a^{2}\cdot st)=s^{8}+a^{4}\cdot t^{8}-is^{4}t^{4}-ia^{4}s^{4}t^{4}=s^{8}-i\left(a^{4}+1\right)\cdot s^{4}t^{4}+a^{4}\cdot t^{8}.$
This is a homogeneous polynomial of degree two in the variables $(s^{4},t^{4})$. Hence, for either one of the eight solutions of the discriminant
(6) $(a^{4}+1)^{2}+4a^{4}=a^{8}+6a^{4}+1=0$
we find two components $\{C_{j,1},C_{j,2}\}_{j=1}^{8}$ of $\Phi^{-1}(\Delta_{a})$. These are naturally sections of either one of the two fibrations $\pi_{i}$. In order to gain computational access to such solutions for $a$, we extend the field over which we are working from $\mathbb{Q}(\zeta_{8})$ to the splitting field $K$ of the above discriminant. The degree of the extension is $[K:\mathbb{Q}(\zeta_{8})]=2$.
These computations are carried out in lines 75-97 of minimum_entropy_surface.jl. Note that in that script we start out with the field $K$ from the beginning in order to avoid complicated base changes later on.
Remark 4.2.
In this first construction of $X_{h}\cong X_{OY}$, we were also able to extract the components of the reducible fibers and realize the action by the group $G$ on $X_{h}$. As predicted by Proposition 3.11 the sublattice $L$ of $\mathrm{NS}(X_{h})$ spanned by the curves of degree $2$ we found has rank $18$ and determinant $1024=[\mathrm{NS}(X_{h}):L]\det(\mathrm{NS}(X_{h}))$.
4.3. Extracting a Weierstrass model
The construction of $X_{h}\overset{2:1}{\longrightarrow}\mathbb{P}^{1}\times\mathbb{P}^{1}\to\mathbb{P}^{1}$ as a $2:1$-cover already suggests an entry point for the computation of a Weierstrass equation. Passing to the affine chart $U$ where $x=1$ of the ambient space $\mathbb{P}(\mathcal{O}\oplus\mathcal{O}(-2)|_{Q})$ of $X_{h}$ with coordinates $(\lambda,y,z,w)$, we can use the defining equation $q=xy+zw$ for $Q$ to eliminate $y$. Then we are left with
$\lambda^{2}=1+z^{4}w^{4}-iz^{4}-iw^{4},$
which is the pullback of Equation 4 to the respective affine chart over $\mathbb{P}^{1}\times\mathbb{P}^{1}$ discussed earlier. This is already an equation of a suitable format and classical formulas apply to transform it to a Weierstrass equation; see e.g. the work by A. Kumar [Kumar14] for such procedures. To this end we need to specify a section, i.e. a $K(t)$-rational point in the associated elliptic curve. We choose this section to be one of those constructed earlier.
The half-automated extraction of a globally minimal Weierstrass equation for an elliptic curve over $K(t)$, its simplification and the rational maps for the various identifications are carried out in lines 100-200 of minimum_entropy_surface.jl. We end up with the Weierstrass equation
(7) $y^{2}=x^{3}-(t^{8}+1)x.$
Remark 4.3.
In hindsight, one could have arrived at this equation starting from Proposition 3.9 which gives the existence of an elliptic fibration on $X_{OY}$ with $8$ fibers of type $A_{1}$ and a $2$-torsion section. Since there is a non-symplectic automorphism $\alpha$ of order $8$ preserving the fibers and the section, this in fact determines the Weierstrass equation uniquely: Looking at the action of $\alpha$ on $\mathrm{NS}(X_{OY})$ one sees that $\alpha$ acts with order $8$ on the the base $\mathbb{P}^{1}$. Up to a change of coordinates, $\alpha$ thus acts by $\zeta_{8}$ on $t$. Since there is a $2$-torsion section the Weierstrass equation is of the form $y^{2}=x^{3}+c_{4}(t)x$ with $c_{4}\in K[t]$ of degree $8$, with $8$ simple roots giving the $A_{1}$-singular fibers and invariant under $\alpha$. Hence $c_{4}(t)=at^{8}+b$. After a change of coordinates one may assume $a=b=-1$.
In order to be able to work with smaller extensions of $\mathbb{Q}$, we pass from the equation (7) to the elliptic surface with generic fiber $E_{1}$ defined by
(8) $y^{2}=x^{3}+(1-t^{8})x,$
which is isomorphic to the previous one, as they are related by a change of Weierstrass coordinates
$(x,y,t)\mapsto(-\zeta_{16}^{4}\cdot x,\;-\zeta_{16}^{2}\cdot y,\;\zeta_{16}\cdot t)$
where $\zeta_{16}$ is a primitive $16$-th root of unity. The associated elliptically fibered K3 surface $X_{e_{1}}$ is defined over $\mathbb{Q}$ and its trivial lattice over $\mathbb{Q}(\zeta_{8})$. However, in order to realize all algebraic cycles for the Néron-Severi lattice $\mathrm{NS}(X_{e_{1}})$, we still need the extension $K/\mathbb{Q}(\zeta_{8})$ of degree two generated by the algebraic element $\alpha$ with minimal polynomial given in Equation 6.
4.4. Specialisation to positive characteristic
In practice, it is computationally challenging and thus inconvenient to work over the field of rational numbers and even more so over a number field of degree $8$. Therefore we specialised the equations to positive characteristic when experimenting. The general theory is as follows:
Let $X_{K}$ be K3 surface over a number field $K$ and let $\mathfrak{p}$ be a finite place of $K$ where $X_{K}$ has good reduction. Choose a smooth model of $X_{K}$ over the localized ring of integers of $\mathcal{O}_{K,\mathfrak{p}}$ and let $X_{\mathfrak{p}}$ be its special fiber. Let $X$ denote the base change of $X_{K}$ to the algebraic closure $\overline{K}$ and $X_{\overline{k}}$ the base change of $X_{\mathfrak{p}}$ to the algebraic closure of the residue field $k$. Specialisation of (effective) divisors induces a specialisation map
(9) $\mathrm{sp}:\mathrm{NS}(X)\to\mathrm{NS}(X_{\overline{k}}),$
which is injective [luijk, Proposition 6.2] and preserves intersection products [Fulton98, Corollary 20.3]. Note that in general it fails to be surjective [Charles14] [CostaElsenhansJanel]. Under the additional hypothesis that $\mathrm{sp}$ is surjective and $\mathrm{sp}(\mathcal{L})$ is ample for at least one ample line bundle $\mathcal{L}$ on $X$, the specialisation map $\mathrm{sp}$ is equivariant with respect to the group homomorphism $\mathrm{Aut}(X)\to\mathrm{Aut}(X_{\overline{k}})$ [lieblich-maulik, Theorem 2.1] and the latter is injective.
Let us return to our surface $X_{e_{1}}/\mathbb{Q}$ constructed in the previous section. The choice of the Weierstrass equation $y^{2}=x^{3}+(1-t^{8})x$ defines a model of $X_{e_{1}}$ over $\mathrm{Spec}\,\mathbb{Z}$ which has good reduction away from $2$. Recall that to access $\mathrm{NS}(X_{e_{1}})$, we had to perform a base change to the number field $K/\mathbb{Q}$ of degree $8$. In order to compute over a prime field, one has to choose an odd prime $\mathfrak{p}\leq\mathcal{O}_{K}$ of degree $1$. We picked a prime over $113$ ($41$ works just as well), so that the residue field $k$ equals $\mathbb{F}_{113}$. Finally, note that if $f$ is an automorphism of $X_{e_{1}}/\mathcal{O}_{K}$ and $\overline{f}\in\mathrm{Aut}(X_{k})$ its specialisation, then $\mathrm{sp}\circ f^{*}=\overline{f}^{*}\circ\mathrm{sp}$. These considerations pay off in practice by the fact that both, the correctness of the final result, as well as some intermediate computations of intersection numbers, can be certified in positive characteristic.
For our practical computations, however, it is not necessary to define a model of $E_{1}$ over $\mathbb{Z}$, a ring of algebraic integers, or a localization thereof as a data structure of its own. Instead, we will be working with $E_{1}$ over the number field $K$ introduced above and utilize the projection map
(10) $K\dashrightarrow\mathbb{F}_{113},\quad\frac{p}{q}\mapsto(p+\mathfrak{p})\cdot(q+\mathfrak{p})^{-1}\quad\textnormal{for}\quad q\notin\mathfrak{p}$
which is defined only on the localization $\mathcal{O}_{K,\mathfrak{p}}\subset K$. Should this map ever be applied to a non-admissible element, an error will be thrown to indicate that the curve or scheme’s concrete model is not defined over the ring $\mathcal{O}_{K,\mathfrak{p}}$.
From an experimental point of view, it is much more convenient to continue with $\overline{X}_{e_{1}}$ and search for the specialization $\overline{f}_{OY}\colon(X_{e_{1}})_{k}\to(X_{e_{1}})_{k}$ of the automorphism $f_{OY}$ in characteristic zero. This is what we did in practice. After we found an automorphism $\overline{f}_{OY}$ with the anticipated action on $\mathrm{NS}((X_{e_{1}})_{k})$, we let it guide our search for $f_{OY}$ and matched the objects in characteristic zero to their reductions over $k=\mathbb{F}_{113}$. Only this final calculation is visible and carried out in minimum_entropy_automorphism.jl.
4.5. Fixing generators of the Néron-Severi lattice
The trivial lattice of $X_{e_{1}}$ is automatically extracted in OSCAR from the resolution of singularities of the Weierstrass model. In general, the computation of the Mordell-Weil group of an elliptic curve is difficult and there is no automatism in OSCAR. However, by Remark 4.2, we have found generators of a sublattice $L$ of index $2$ in $\mathrm{NS}(X_{e_{1}})$, consisting of fiber components and sections. These can be transferred to the Weierstrass model. We did the pushforward in positive characteristic. To also find the sections in characteristic zero, i.e. the elements of the Mordell-Weil group, we apply a multivariate Newton iteration to $\mathbb{F}_{113}(t)$-rational points on the reduction $\overline{E}_{1}$ over $\mathbb{F}_{113}(t)$ to obtain $\mathbb{Q}_{113}(t)$-rational points of $E_{1}$ (up to finite precision in the $113$-adic numbers $\mathbb{Q}_{113}$). These are then used to reconstruct $K(t)$-rational points on $E_{1}$ via lattice reduction. Note that since the map $\mathrm{sp}$ in Equation 9 is injective, and the curves considered are unique in their numerical equivalence class, the lift is unique.
The Mordell-Weil group of $\overline{E}_{1}$ is generated by the points with coordinates $(x,y)$ given by
(11) $\begin{matrix}\overline{S_{1}}&=&(62t^{4}+6t^{3}+106t^{2}+75t+26,&66t^{6}+67t^{5}+37t^{4}+60t^{3}+67t^{2}+101t+58)\\ \overline{S_{2}}&=&(112t^{4}+97t^{2}+98,&104t^{5}+82t^{3}+91t)\\ \overline{S_{3}}&=&(t^{4}+112,&87t^{4}+26)\\ \overline{S_{4}}&=&(8t^{4}+17t^{3}+11t^{2}+70t+7,&39t^{6}+17t^{5}+60t^{4}+64t^{3}+41t^{2}+29t+89)\\ \overline{S_{5}}&=&(50t^{4}+104t^{2}+72,&56t^{6}+30t^{4}+36t^{2}+104)\\ \overline{S_{6}}&=&(31t^{4}+45t^{3}+60t^{2}+59t+13,&43t^{6}+18t^{5}+67t^{4}+75t^{3}+10t^{2}+44t+17)\\ \overline{S_{7}}&=&(105t^{4}+43t^{3}+61t^{2}+17t+7,&20t^{6}+33t^{5}+53t^{4}+9t^{3}+63t^{2}+70t+89)\\ \overline{S_{8}}&=&(62t^{4}+75t^{3}+8t^{2}+107t+87,&66t^{6}+103t^{5}+103t^{4}+63t^{3}+46t^{2}+37t+34)\\ \end{matrix}$
Liftings $S_{i}$ of these points to $E_{1}/K(t)$ are given in Figure 1. We let $s_{i}=[S_{i}]\in\mathrm{NS}(X_{e_{1}})$ for $i=1,\dots,8$.
Let $o_{1}$ be the zero section, and $e_{1}$ the class of a fiber. By Tate’s algorithm [Silverman94, IV §9 Table 4.1], we have eight reducible fibers at $t^{8}=1$ consisting of two $(-2)$ curves meeting tangentially in one point. These are said to be of Kodaira type III and form and $\widetilde{A}_{1}$-configuration. Each one of these fibers provides us with a single component $a_{1,i}$ which does not meet $o_{1}$. We fix the indices $i=1,\dots,8$ in such a way that $a_{1,1},\dots a_{8,1}$ are contained in fibers over the points
$\begin{array}[]{lccclcc}t_{1}&=&-1,&&t_{2}&=&-\frac{1}{4}\alpha^{6}-\frac{1}{4}\alpha^{4}-\frac{7}{4}\alpha^{2}-\frac{3}{4},\\ t_{3}&=&\frac{1}{4}\alpha^{6}-\frac{1}{4}\alpha^{4}+\frac{7}{4}\alpha^{2}-\frac{3}{4},&&t_{4}&=&-\frac{1}{2}\alpha^{6}-\frac{5}{2}\alpha^{2},\\ t_{5}&=&\frac{1}{2}\alpha^{6}+\frac{5}{2}\alpha^{2},&&t_{6}&=&-\frac{1}{4}\alpha^{6}+\frac{1}{4}\alpha^{4}-\frac{7}{4}\alpha^{2}+\frac{3}{4},\\ t_{7}&=&\frac{1}{4}\alpha^{6}+\frac{1}{4}\alpha^{4}+\frac{7}{4}\alpha^{2}+\frac{3}{4},&&t_{8}&=&1.\end{array}$
where $\alpha$ is fixed with $\alpha^{8}+6\alpha^{4}+1=0$.
Let $B=(e_{1},o_{1},a_{1,1},\dots a_{1,8},s_{1},\dots s_{8})$. The gram matrix with respect to the basis $B$ of $\mathrm{NS}(X_{e_{1}})\otimes\mathbb{Q}$ is
$G_{B}=\left(\begin{array}[]{rrrrrrrrrrrrrrrrrr}0&1&0&0&0&0&0&0&0&0&1&1&1&1&1&1&1&1\\ 1&-2&0&0&0&0&0&0&0&0&0&0&0&0&0&0&0&0\\ 0&0&-2&0&0&0&0&0&0&0&1&0&1&0&1&1&1&0\\ 0&0&0&-2&0&0&0&0&0&0&0&1&0&0&0&0&1&1\\ 0&0&0&0&-2&0&0&0&0&0&1&0&0&1&1&0&1&1\\ 0&0&0&0&0&-2&0&0&0&0&1&1&1&0&0&0&1&0\\ 0&0&0&0&0&0&-2&0&0&0&0&1&1&1&0&1&0&1\\ 0&0&0&0&0&0&0&-2&0&0&0&0&0&0&1&1&0&0\\ 0&0&0&0&0&0&0&0&-2&0&1&1&0&1&0&1&0&0\\ 0&0&0&0&0&0&0&0&0&-2&0&0&1&1&1&0&0&1\\ 1&0&1&0&1&1&0&0&1&0&-2&2&2&2&2&0&0&2\\ 1&0&0&1&0&1&1&0&1&0&2&-2&1&0&2&2&1&1\\ 1&0&1&0&0&1&1&0&0&1&2&1&-2&0&0&1&2&1\\ 1&0&0&0&1&0&1&0&1&1&2&0&0&-2&1&2&2&1\\ 1&0&1&0&1&0&0&1&0&1&2&2&0&1&-2&1&1&1\\ 1&0&1&0&0&0&1&1&1&0&0&2&1&2&1&-2&2&2\\ 1&0&1&1&1&1&0&0&0&0&0&1&2&2&1&2&-2&1\\ 1&0&0&1&1&0&1&0&0&1&2&1&1&1&1&2&1&-2\end{array}\right).$
The lattice $\mathrm{NS}(X_{e_{1}})$ is generated by $B$ together with the $2$-torsion section $p_{2}$. The latter has the following coordinates with respect to $B$.
$v_{B}(p_{2})=\frac{1}{2}\left(\begin{array}[]{rrrrrrrrrrrrrrrrrr}4&2&-1&-1&-1&-1&-1&-1&-1&-1&0&0&0&0&0&0&0&0\end{array}\right).$
The $G$-invariant ample class is given by
$v_{B}(h)=\frac{1}{2}\left(\begin{array}[]{rrrrrrrrrrrrrrrrrr}8&2&-1&-1&-1&1&-1&-1&1&-1&4&4&0&-2&2&-2&-4&0\end{array}\right).$
Let $\epsilon$ denote the involution on $X_{e_{1}}$ given in Weierstrass coodinates by $(x,y,t)\mapsto(x,y,-t)$. By direct computations we find that the Enriques involution $j\in G$ of $X_{h}$ is given on $X_{e_{1}}$ by $T_{P}\circ\epsilon$ where $T_{P}$ denotes the translation automorphism induced by the section $P$ with coordinates
$x$ $=$ $\frac{(-\alpha^{6}-\alpha^{4}-7\alpha^{2}+1)t+4}{4t+\alpha^{6}-\alpha^{4}+7\alpha^{2}+1},$
$y$ $=$ $\frac{2}{2t^{3}+(\alpha^{6}-\alpha^{4}+7\alpha^{2}+1)t^{2}+(-\alpha^{4}+\alpha^{2}-1)t}\cdot\left(\left(\frac{3}{2}\alpha^{7}-\alpha^{5}+\frac{19}{2}\alpha^{3}-8\alpha\right)t^{4}\right.$
$+\left(-10\alpha^{7}-2\alpha^{5}-58\alpha^{3}-18\alpha\right)t^{3}+\left(-12\alpha^{7}+6\alpha^{5}-72\alpha^{3}+30\alpha\right)t^{2}$
$\left.+\left(8\alpha^{7}+4\alpha^{5}+44\alpha^{3}+24\alpha\right)t+\frac{7}{2}\alpha^{7}-\alpha^{5}+\frac{39}{2}\alpha^{3}-4\alpha\right).$
To obtain a simpler Enriques involution on $X_{e_{1}}$ compose $\epsilon$ with the translation by the unique $2$-torsion section. The resulting involution $\iota$ is an Enriques involution and $X/\iota$ is an Enriques surface of base-change type. Using our OSCAR implementation, we compute the coinvariant lattice $\mathrm{NS}(X_{e_{1}})_{\iota}$ and check that it is in the correct genus. Then Proposition 3.5 yields that $\iota$ is $\mathrm{Aut}(Y)$-conjugate to $\iota_{OY}$ and that $X_{e_{1}}/\iota:=Y_{f_{1}}\cong Y_{OY}$.
It remains to express $(f_{OY})_{*}$ and $(\iota_{OY})_{*}$ in terms of the basis $B$ of $\mathrm{NS}(X_{e_{1}})$. So far we have it only as an isometry on the lattice $L_{K3}$ coming from the database entry 40.8.0.1. It comes with an implicitly given Hodge structure and $G$-invariant ample class $h$.
We thus have to define an embedding $\gamma\colon\mathrm{NS}(X_{h})\to L_{K3}$ that can be extended to a Hodge isometry $H^{2}(X_{h},\mathbb{Z})\to L_{K3}$. Recall from Proposition 3.11 that in $L_{K3}$ we are given $64=32+32$ vectors $(c_{1},\dots,c_{64})=C$ of square $(-2)$. They correspond to $64$ explicit classes $D=(d_{1},\dots,d_{64})$ spanning an index $2$ sublattice of $\mathrm{NS}(X_{h})$. One can view $C$ and $D$ as graphs with edge-weights, compute a graph isomorphism [^3] We used sageMath [sagemath] for this purpose. $\delta\colon C\to D$ and define $\gamma$ as its linear extension (if it exists). Then, one has to modify the isomorphism in such a way that it extends to a Hodge isometry. Finally, one computes the pushforward $\mathrm{NS}(X_{h})\to\mathrm{NS}(X_{e_{1}})$, to represent everything with respect to the basis $B$ of the elliptic fibration on $X_{e_{1}}$.
A more direct way to go forward is the following: We compute the image of $\mathrm{Aut}(Y_{f_{1}})\to O(\mathrm{Num}(Y_{f_{1}}))$ and use it to find an automorphism of the correct entropy. Since the intermediate results and details of these computations are not of relevance, we just present the end result and how it can be checked for correctness.
The matrix representation of the pushforward $(f_{OY})_{*}$ on $\mathrm{NS}(X_{e_{1}})$ with respect to the basis $B$ is
(12) $D_{B}((f_{OY})_{*})=\left(\begin{array}[]{cccccccccccccccccc}3&1&-1&-1&1&-1&0&-1&1&0&1&0&-1&1&0&0&0&0\\ 2&1&0&0&0&0&-1&-1&0&0&1&0&0&0&0&-1&0&0\\ 3&1&-1&-1&1&-1&-1&-1&1&0&1&0&-1&1&0&0&0&0\\ 2&1&-1&0&1&-1&0&-1&1&0&1&0&-1&1&0&0&0&0\\ 3&1&-1&-1&0&-1&0&-1&1&0&1&0&-1&1&0&0&0&0\\ 2&1&0&-1&1&-1&0&-1&1&0&1&0&-1&1&0&0&0&0\\ 0&0&0&0&0&0&0&0&0&1&0&0&0&0&0&0&0&0\\ 1&0&0&0&0&0&0&-1&0&0&0&0&0&0&0&0&0&0\\ 0&0&0&0&0&0&0&0&1&0&0&0&0&0&0&0&0&0\\ 1&0&0&0&0&-1&0&0&0&0&0&0&0&0&0&0&0&0\\ -3&-2&\frac{1}{2}&\frac{1}{2}&\frac{3}{2}&\frac{1}{2}&\frac{1}{2}&-\frac{1}{2}&\frac{1}{2}&\frac{1}{2}&1&0&0&1&0&0&1&0\\ 14&6&-\frac{5}{2}&-\frac{5}{2}&-\frac{1}{2}&-\frac{3}{2}&-\frac{5}{2}&-\frac{5}{2}&-\frac{1}{2}&-\frac{3}{2}&2&0&-1&0&0&-2&-1&-1\\ 8&4&-2&-1&0&-1&-1&-2&0&-1&1&0&-1&0&0&-1&0&0\\ 6&3&-\frac{3}{2}&-\frac{3}{2}&\frac{3}{2}&-\frac{3}{2}&-\frac{3}{2}&-\frac{3}{2}&\frac{1}{2}&-\frac{1}{2}&2&0&-2&1&1&-1&-1&0\\ 10&5&-2&-1&0&-1&-1&-2&0&-1&1&0&-1&0&-1&-1&0&0\\ 2&1&-\frac{1}{2}&-\frac{1}{2}&\frac{1}{2}&-\frac{1}{2}&\frac{1}{2}&-\frac{3}{2}&\frac{1}{2}&-\frac{1}{2}&1&0&0&1&-1&0&1&0\\ 0&-1&0&0&1&1&0&-1&1&0&2&1&0&0&0&-1&0&0\\ 10&4&-\frac{5}{2}&-\frac{5}{2}&\frac{5}{2}&-\frac{3}{2}&-\frac{3}{2}&-\frac{7}{2}&\frac{3}{2}&-\frac{1}{2}&3&0&-2&2&0&-2&0&0\end{array}\right)$
and the pushforward of the accompanying Enriques involution reads
(13) $D_{B}(\iota_{*})=\left(\begin{array}[]{cccccccccccccccccc}1&0&0&0&0&0&0&0&0&0&0&0&0&0&0&0&0&0\\ 2&1&-\frac{1}{2}&-\frac{1}{2}&-\frac{1}{2}&-\frac{1}{2}&-\frac{1}{2}&-\frac{1}{2}&-\frac{1}{2}&-\frac{1}{2}&0&0&0&0&0&0&0&0\\ 1&0&0&0&0&0&0&0&0&-1&0&0&0&0&0&0&0&0\\ 1&0&0&0&0&0&0&0&-1&0&0&0&0&0&0&0&0&0\\ 1&0&0&0&0&0&0&-1&0&0&0&0&0&0&0&0&0&0\\ 1&0&0&0&0&0&-1&0&0&0&0&0&0&0&0&0&0&0\\ 1&0&0&0&0&-1&0&0&0&0&0&0&0&0&0&0&0&0\\ 1&0&0&0&-1&0&0&0&0&0&0&0&0&0&0&0&0&0\\ 1&0&0&-1&0&0&0&0&0&0&0&0&0&0&0&0&0&0\\ 1&0&-1&0&0&0&0&0&0&0&0&0&0&0&0&0&0&0\\ 2&1&-1&0&0&-1&0&0&0&0&0&0&-1&1&0&0&0&0\\ 4&2&-\frac{1}{2}&-\frac{1}{2}&-\frac{1}{2}&-\frac{1}{2}&-\frac{1}{2}&-\frac{1}{2}&-\frac{1}{2}&-\frac{1}{2}&0&-1&0&0&0&0&0&0\\ 0&0&\frac{1}{2}&-\frac{1}{2}&-\frac{1}{2}&\frac{1}{2}&\frac{1}{2}&-\frac{1}{2}&-\frac{1}{2}&\frac{1}{2}&0&0&1&0&0&0&0&0\\ -2&-1&1&0&0&1&0&0&0&0&1&0&1&0&0&0&0&0\\ 0&0&\frac{1}{2}&-\frac{1}{2}&\frac{1}{2}&-\frac{1}{2}&-\frac{1}{2}&\frac{1}{2}&-\frac{1}{2}&\frac{1}{2}&0&0&0&0&1&0&0&0\\ 2&1&-1&0&0&-1&0&0&0&0&-1&0&-1&1&0&1&0&0\\ 4&2&-1&-1&0&-1&-1&0&0&0&1&0&-1&0&1&-1&-1&0\\ 12&6&-1&-1&-3&0&-2&-1&-2&-2&-1&0&1&-2&-1&-1&0&-1\end{array}\right).$
At this point one should perform a number of sanity checks on the matrix $D_{B}((f_{OY})_{*})$:
  1. (1)
    it is an isometry, i.e. $D_{B}((f_{OY})_{*})G_{B}D_{B}((f_{OY})_{*})^{T}=G_{B}$
  2. (2)
    its spectral radius is $\tau_{8}$,
  3. (3)
    it preserves the nef cone, i.e. $h^{\prime}:=(f_{OY})_{*}(h)$ is ample,
  4. (4)
    it can be extended to a Hodge isometry,
  5. (5)
    it commutes with $D_{B}(\iota_{*})$.
(1) and (2) are trivial to check using any computer algebra system. For (3) one certifies that the set
$\{r\in\mathrm{NS}(X_{e_{1}})\mid r^{2}=-2,r.h>0,r.h^{\prime}<0\}$
is empty. It is computed using the OSCAR function separating_hyperplanes based on [Shimada24, §3]. For (4) one computes its action on the discriminant group and checks that it generates the same subgroup of $O(\mathrm{NS}(X)^{\vee}/\mathrm{NS}(X))$, as the action of the non-symplectic automorphism of order $8$ given by $t\mapsto\zeta_{8}t$. These checks are carried out in the file sanity_check.jl.
5. Construction of the automorphism
In order to find $f_{OY}$, we apply techniques called “Kneser’s neighbor method” and “fibration hopping”. These have been described in [Kumar14] and [ElkiesKumar14] and they have already been put to use in [ElkiesSchuett15], [Kumar15], and [BrandhorstElkies23]. As it turned out, we do not need to apply these methods in full generality to solve our particular problem and we shall henceforth confine ourselves to describing the minimal requirements only.
Given that we already know the prescribed action of $f_{OY}$ on the Néron-Severi lattice $\mathrm{NS}(X_{e_{1}})$ of our K3 surface, Equation 12, we may consider the image of the class of a fiber of $\pi_{1}$
$e_{1}^{\prime}:=(f_{OY})_{*}(e_{1})\in\mathrm{NS}(X_{e_{1}}).$
Its basis representation is given by the first row of the matrix in Equation 12. Since $f_{OY}$ is an automorphism of the very same surface, $e^{\prime}$ appears as the class of a fiber of another elliptic fibration
$\pi^{\prime}\colon X_{e_{1}^{\prime}}\to\mathbb{P}^{1}$
which must be isomorphic to the original one by means of $f_{OY}$. Our next goal is to explicitly compute the elliptic fibration $\pi^{\prime}$. This can be done by fibration hopping, which in a nutshell is the following.
Recall that elliptic fibrations on a K3 surface $X$ are in bijection with primitive, isotropic and nef classes $e\in\mathrm{NS}(X)$. Such classes are called elliptic. Indeed the linear system $|e|$ of an elliptic class induces an elliptic fibration (possibly without section) and conversely, the class of a fiber of any elliptic fibration is an elliptic class. Two elliptic fibrations corresponding to elliptic classes $e,e^{\prime}\in\mathrm{NS}(X)$ are called $n$-neighbors if $e.e^{\prime}=n$.
Given an elliptic fibration $\phi\colon X\to\mathbb{P}^{1}$ with elliptic class $e$ and an elliptic class $e^{\prime}\in\mathrm{NS}(X)$ with $e.e^{\prime}=2$, i.e. the class of a $2$-neighborly fibration, a two neighbor step is a procedure to compute the fibration $\phi^{\prime}\colon X\to\mathbb{P}^{1}$ corresponding to $e^{\prime}$ in terms of a weierstrass model and the corresponding changes of coordinates.
In our case $e_{1}.e_{1}^{\prime}=2$ and so $\phi:=\pi_{1}$ and $\phi^{\prime}:=\pi^{\prime}$ are $2$-neighbors.
5.1. A $2$-neighbor step
We proceed as follows. First we compute the linear system of $e^{\prime}$ by means of the methods outlined in [BrandhorstElkies23, Lemma 2.14]. This provides us with an elliptic-fibration $\pi^{\prime}\colon X\to\mathbb{P}^{1}$, described by a rational function on $X_{e_{1}}$, the pullback of an affine coordinate of $\mathbb{P}^{1}$.
A well established procedure (which has, for example, also been described and used in [BrandhorstElkies23]) allows to extract a new equation of the form
$\tilde{y}^{2}=g(\tilde{x}),\quad g(\tilde{x})\in K(\tilde{t})[\tilde{x}]\textnormal{ of degree }\leq 4,$
from the original Weierstrass equation (8), together with the rational change of coordinates
$\Phi\colon K(\tilde{t},\tilde{x},\tilde{y})\to K(t,x,y)$
taking (8) to $\tilde{y}^{2}-g(\tilde{x})$. This has been automatized in the function two_neighbor_step called in
minimum_entropy_automorphism.jl.
Next, we need to find a section for $\pi^{\prime}$, i.e. a $K(\tilde{t})$-rational point. Finding rational points is known to be a hard problem. Hence, it still has to be carried out by hand and constitutes the main obstruction to full automatization of $2$-neighbor steps.
We are lucky since the section $[D]:=s_{8}$ of $\pi$ satisfies $s_{8}.e^{\prime}=1$. Hence its pushforward $\Phi_{*}D$ induces a $K(\tilde{t})$-rational point. However, computations for the extraction of explicit coordinates as a point on the curve defined by $\tilde{y}^{2}-g(\tilde{x})$ do not terminate in OSCAR at the moment. For the reduction in positive characteristic computations do terminate in OSCAR and we can match the result with the reduction of the hardcoded point in characteristic zero.
This information, in turn, provides the input to a constructor for an elliptic surface which automatically takes care of transforming the given equation into a globally minimal integral Weierstrass form and returns the elliptic surface, together with the additional rational transformation. The final Weierstrass equation defining our second model $X_{e^{\prime}}$ for this elliptic surface is
$\tilde{y}^{2}$ $=$ $\tilde{x}^{3}-((963\alpha^{7}+468\alpha^{6}+27\alpha^{5}-168\alpha^{4}+5613\alpha^{3}+2724\alpha^{2}+165\alpha-987)\tilde{t}^{8}$
$+(8146\alpha^{7}+8200\alpha^{6}+4034\alpha^{5}+328\alpha^{4}+47486\alpha^{3}+47720\alpha^{2}+23654\alpha+1776)\tilde{t}^{7}$
$+(-8946\alpha^{7}+29456\alpha^{6}+30142\alpha^{5}+15750\alpha^{4}-52094\alpha^{3}+171080\alpha^{2}+176890\alpha+90594)\tilde{t}^{6}$
$+(-198380\alpha^{7}-23604\alpha^{6}+59304\alpha^{5}+66080\alpha^{4}-1156092\alpha^{3}-140588\alpha^{2}+351904\alpha+378672)\tilde{t}^{5}$
$+(-505540\alpha^{7}-269710\alpha^{6}-41160\alpha^{5}+81760\alpha^{4}-2946020\alpha^{3}-1581790\alpha^{2}-219520\alpha+455420)\tilde{t}^{4}$
$+(-376040\alpha^{7}-426636\alpha^{6}-243656\alpha^{5}-29148\alpha^{4}-2190552\alpha^{3}-2506924\alpha^{2}-1378328\alpha-212716)\tilde{t}^{3}$
$+(129024\alpha^{7}-187768\alpha^{6}-238056\alpha^{5}-118762\alpha^{4}+753592\alpha^{3}-1120308\alpha^{2}-1334368\alpha-746354)\tilde{t}^{2}$
$+(232488\alpha^{7}+47498\alpha^{6}-62932\alpha^{5}-62418\alpha^{4}+1356168\alpha^{3}+258014\alpha^{2}-328156\alpha-403254)\tilde{t}$
$+53663\alpha^{7}+32688\alpha^{6}+4351\alpha^{5}-4928\alpha^{4}+313121\alpha^{3}+184520\alpha^{2}+37697\alpha-41349)\cdot\tilde{x}$
The rational transformations produced by this process can be used to construct an explicit identification
$\Xi\colon X_{e^{\prime}}\overset{\cong}{\longrightarrow}X_{e_{1}}$
of our two instances of the K3 surface arising from the different Weierstrass models. This is what we call a geometric realization of a $2$-neighbor step.
Remark 5.1.
Note that the identification $\Xi$ is not unique! It can, for instance be modified by the automorphism of $X_{e^{\prime}}$ given by translation with a section, or the one induced by an admissible Möbius transformation (see Definition 5.2) on the base $\mathbb{P}^{1}$ of the elliptic fibration. In fact, this ambiguity will become important below.
Moreover, these ambiguities and the related automated choices introduce a certain randomness in the intermediate outputs of our construction script minimum_entropy_automorphism.jl. It is important to note that therefore many of the following matrices and other intermediate results may differ in a reproduction of our findings; we nevertheless leave them here for illustration and comparison.
5.2. Completion to an automorphism of $X_{e_{1}}$
Up to this point we produced the identification $\Xi\colon X_{e^{\prime}}\to X_{e_{1}}$. Since the elliptic fibrations $\pi\colon X_{e_{1}}\to\mathbb{P}^{1}$ and $\pi^{\prime}\colon X_{e^{\prime}}\to\mathbb{P}^{1}$ are abstractly isomorphic, we can, on the one hand, think of $h$ as a “change of (Weierstrass-) coordinates” on the same surface. On the other hand, we may choose an identification $\Psi$ of these two different descriptions of this fibration.
[…]
The composition $\Xi\circ\Psi\colon X_{e_{1}}\to X_{e_{1}}$ is then already an honest automorphism of $X_{e_{1}}$; the latter also considered as a single instance of this object in memory.
To get our hands on possible choices for $\Psi$, we implemented the internal function
admissible_moebius_transformations which takes two EllipticSurfaces as input.
Definition 5.2.
Given two elliptic fibrations $\pi,\pi^{\prime}\colon X\to\mathbb{P}^{1}$, a Möbius transformation $\mu\colon\mathbb{P}^{1}\to\mathbb{P}^{1}$ is admissible if it lifts to an automorphism $\Psi\colon X\to X$ such that $\pi^{\prime}\circ\Psi=\mu\circ\pi$.
Consequently, the function computes a full set of those Möbius transformations of the base $\mathbb{P}^{1}$ which take the configuration of critical values of the first elliptic fibration into the configuration of the second. Note that this list might already be empty, but it is certainly finite once we have at least three such critical values.
In a second step, an attempt is made to lift every such transformation $\mu\colon\mathbb{P}^{1}\to\mathbb{P}^{1}$ to an isomorphism of the elliptic surfaces. Whether or not such an identification exists is decided from the generic fibers after base change by $\mu$ and in the affirmative case such an isomorphism is automatically extended to the associated elliptic surfaces.
For such an arbitrary choice of $\Psi$ we then compute the pushforward of $\Xi\circ\Psi$ on $\mathrm{NS}(X_{e_{1}})$. This is done in a highly automated and optimized way via the method pushforward_on_algebraic_lattices, but uses parts of the generic machinery developed for algebraic cycles; see the next section on implementations for more details.
The explicit form of $(\Xi\circ\Psi)_{*}$ then gives us a first impression on how far we still are from $f_{OY}$. Note that, by construction, $\Xi\circ\Psi$ is an automorphism of $X_{e_{1}}$ which takes the class of the fiber $F$ to the class of $F^{\prime}=(f_{OY})_{*}(F)$, but $\Xi\circ\Psi$ and $f_{OY}$ will in general be different at this point.
5.3. Resolving the final ambiguity
To get from $\Xi\circ\Psi$ to $f_{OY}$ still requires some massaging. First we may browse through the lifts of admissible Möbius automorphisms for $X_{e_{1}}$ and select one, $\Phi$ over $\nu$ say, for which $\Xi\circ\Psi\circ\Phi$ takes the reducible fibers of $X_{e_{1}}$ to the same ones as $f_{OY}$.
[…]
Whether or not this is the case can be read off from the matrices for the respective pushforwards on $\mathrm{NS}(X_{e_{1}})$. To this end we may inspect the pushforward of the composite map $(f_{OY}^{-1})_{*}\circ(\Xi\circ\Psi)_{*}\circ\Phi_{*}$ on $\mathrm{NS}(X_{e_{1}})$, which we can build from the already known information on the Hodge and lattice theoretic side.
Once we have found and chosen $\Phi$, the composite map $c=f_{OY}^{-1}\circ\Xi\circ\Psi\circ\Phi$ is an honest automorphism of the elliptic fibration $\pi\colon X_{e_{1}}\to\mathbb{P}^{1}$, i.e. $\pi=\pi\circ c$. At this point the representing matrix for the pushforward of this composite map $c$ on $\mathrm{NS}(X_{e_{1}})$ for the set of generators chosen earlier reads444cf. Remark 5.1
$\begin{pmatrix}\begin{array}[]{c|c|cccccccc|cccccccc}1&0&0&0&0&0&0&0&0&0&0&0&0&0&0&0&0&0\\ \hline\cr-8&-4&\frac{1}{2}&\frac{1}{2}&\frac{5}{2}&-\frac{1}{2}&\frac{3}{2}&\frac{1}{2}&\frac{3}{2}&\frac{3}{2}&1&0&-1&2&1&1&0&1\\ \hline\cr 1&0&-1&0&0&0&0&0&0&0&0&0&0&0&0&0&0&0\\ 0&0&0&1&0&0&0&0&0&0&0&0&0&0&0&0&0&0\\ 0&0&0&0&1&0&0&0&0&0&0&0&0&0&0&0&0&0\\ 1&0&0&0&0&-1&0&0&0&0&0&0&0&0&0&0&0&0\\ 0&0&0&0&0&0&1&0&0&0&0&0&0&0&0&0&0&0\\ 1&0&0&0&0&0&0&-1&0&0&0&0&0&0&0&0&0&0\\ 1&0&0&0&0&0&0&0&-1&0&0&0&0&0&0&0&0&0\\ 0&0&0&0&0&0&0&0&0&1&0&0&0&0&0&0&0&0\\ \hline\cr-6&-3&\frac{1}{2}&\frac{1}{2}&\frac{3}{2}&-\frac{1}{2}&\frac{3}{2}&\frac{1}{2}&\frac{3}{2}&\frac{3}{2}&0&0&-1&2&1&1&0&1\\ -5&-3&\frac{1}{2}&-\frac{1}{2}&\frac{5}{2}&-\frac{1}{2}&\frac{1}{2}&\frac{1}{2}&\frac{3}{2}&\frac{3}{2}&1&-1&-1&2&1&1&0&1\\ -5&-3&\frac{1}{2}&\frac{1}{2}&\frac{5}{2}&-\frac{1}{2}&\frac{1}{2}&\frac{1}{2}&\frac{3}{2}&\frac{1}{2}&1&0&-2&2&1&1&0&1\\ -5&-3&\frac{1}{2}&\frac{1}{2}&\frac{3}{2}&-\frac{1}{2}&\frac{1}{2}&\frac{1}{2}&\frac{3}{2}&\frac{1}{2}&1&0&-1&1&1&1&0&1\\ -5&-3&\frac{1}{2}&\frac{1}{2}&\frac{3}{2}&-\frac{1}{2}&\frac{3}{2}&\frac{1}{2}&\frac{3}{2}&\frac{1}{2}&1&0&-1&2&0&1&0&1\\ -6&-3&\frac{1}{2}&\frac{1}{2}&\frac{5}{2}&-\frac{1}{2}&\frac{1}{2}&\frac{1}{2}&\frac{3}{2}&\frac{3}{2}&1&0&-1&2&1&0&0&1\\ -5&-3&\frac{1}{2}&-\frac{1}{2}&\frac{3}{2}&-\frac{1}{2}&\frac{3}{2}&\frac{1}{2}&\frac{3}{2}&\frac{3}{2}&1&0&-1&2&1&1&-1&1\\ -4&-3&\frac{1}{2}&-\frac{1}{2}&\frac{3}{2}&-\frac{1}{2}&\frac{1}{2}&\frac{1}{2}&\frac{3}{2}&\frac{1}{2}&1&0&-1&2&1&1&0&0\end{array}\end{pmatrix}.$
Recall that with our choices made, the first $10$ generators span the trivial lattice $\mathrm{Triv}(X_{e_{1}})$ while the remaining $8$ ones are sections of the elliptic fibration. Note that the first row of the matrix above is the image of $e_{1}$. We see that $c_{*}(e_{1})=e_{1}$ as expected. For the class of a singular fiber $a_{1,i}+a_{2,i}$, we see in rows $3-10$ that $c_{*}(a_{1,i})$ is either $a_{1,i}$ or $a_{2,i}=e_{1}-a_{1,i}$, confirming that $c_{*}$ acts trivially on $\mathbb{P}^{1}$ (it has $8$ fixed points on $\mathbb{P}^{1}$ below the reducible singular fibers).
The next step to resolve in the ambiguity for the choice of $\Psi$ are translations by sections. The quotient $\mathrm{NS}(X_{e_{1}})/\mathrm{Triv}(X_{e_{1}})$ is isomorphic to the Mordell-Weil group of the generic fiber $E_{1}$ of $X_{e_{1}}$ (cf. [SchuettShioda19, Theorem 6.5]). A closer inspection of the second row of the above matrix, i.e. the image of the zero section $o$ on $X_{e_{1}}$, therefore yields that $o$ is taken to the section $p:=[P]$ with $P$ defined by the linear combination
$P:=P_{2}+S_{1}-S_{3}+2S_{4}+S_{5}+S_{6}+S_{8}$
in the group law of the generic fiber. Here the two torsion section $P_{2}$ comes in because of the denominator $2$ in the second row of the matrix. Therefore, we compose with the translation $T$ by $-P$ on $X_{e_{1}}$.
The very last ambiguity in the choice of $\Psi$ which is left at this point amounts to an automorphism of the generic fiber $E_{1}$. In this case, there are four of them which we can get access to via automorphism_list. Their associated automorphisms of $X_{e_{1}}$ can be computed automatically via the internal function
isomorphism_from_generic_fibers and the induced maps on $\mathrm{NS}(X_{e_{1}})$ as usual with
pushforward_on_algebraic_lattices. It is now a simple matter of trial and error among the four automorphisms to find the correct final compensation $A$. By construction and since the action of $\mathrm{Aut}(X_{OY})$ on the Néron-Severi lattice is faithful (an element in the kernel would act trivially on $\mathrm{NS}(X_{e_{1}})^{\vee}/\mathrm{NS}(X_{e_{1}})\cong T(X_{OY})^{\vee}/T(X_{OY})$ and thus on $T(X_{e_{1}})$), we have constructed the automorphism of Oguiso and Yu as
(16) $\mathrm{id}=A\circ T\circ f_{OY}^{-1}\circ\Xi\circ\Psi\circ\Phi$$\Leftrightarrow$ $f_{OY}=\Xi\circ\Psi\circ\Phi\circ A\circ T.$
Since $(f_{OY})_{*}$ and $j_{*}$ commute, so do $f_{OY}$ and the enriques involution $j$. Thus $f_{OY}$ descends to the Enriques surface $Y_{OY}$.
The final result can be checked for correctness by computing the pushforward $(f_{OY})_{*}$ explicitly, calculating its characteristic polynomial and checking that it divisible by the minimal polynomial $1-x^{2}-2x^{3}-x^{4}+x^{6}$ of $\tau_{8}$.
6. On the implementation
The search for the automorphism of Oguiso and Yu relies heavily on computational tools, the development of which have been pushed in OSCAR with an eye towards this particular application. It is therefore of central interest for this paper to also report on these implementations.
In particular, we have built up a framework for algebraic schemes [^5] See also [BrandhorstZachFruehbis] where we report on work in progress on the framework. throughout the roughly last three years as part of the project B5 of the SFB-TRR 195, Project-ID 286237555, by the DFG. This joint effort has been carried out in wide parts by M. Zach, with significant contributions from S. Brandhorst and A. Frühbis-Krüger and in exchange with J. Böhm. The experience from a previous implementation of resolution of singularities in Singular [Sin] and the efforts for parallelization in algebraic geometry from the joint work of A. Frübis-Krüger and J. Böhm [BoehmFruehbisKrueger18] have been influential for the development.
The following is an overview on the cornerstones of these recent developments in OSCAR which are relevant for this particular publication. The discussion is by no means exhaustive and we make no attempt to prove all details with full mathematical rigor here, as this would expand the paper beyond any form acceptable for publication in a mathematical journal. Instead, we believe that the source code, which is Open Source and available online, can stand for itself, as it can be checked for mathematical rigor and correctness of implementation. For this note we refer to OSCAR version 1.3.0-DEV on commit b4d5f32cbbcaa3bf68fdc798a3c4bc90ace3d758 in the official development branch of OSCAR.
6.1. Polynomial rings, quotients, and localizations
Mathematically, an affine scheme of finite type over a ring $\mathbf{k}$ is of the form $X=\mathrm{Spec}\,R$ for some finitely generated $\mathbf{k}$-algebra $R$. The latter can be a polynomial ring over $\mathbf{k}$, or a quotient thereof. By virtue of Rabinowitsch’s trick this can be extended to localizations
$R[f^{-1}]\cong R[t]/\langle 1-t\cdot f\rangle$
at the multiplicative set given by the powers of an element $f\in R$. This is important as the spectra of such localizations form a basis for the Zariski topology of $\mathrm{Spec}\,R$. Therefore, a first step to realizing (affine) algebraic geometry in OSCAR is to create the data types and functionality for these rings:
  • MPolyRing: multivariate polynomial rings $P$ over a ring $\mathbf{k}$;
  • MPolyQuoRing: quotient rings $R=P/I$ for ideals $I\subset P$;
  • MPolyLocRing: localizations $U^{-1}P$ of polynomial rings $P$ at multiplicative sets $U\subset P$. For the latter we support [^6] We deliberately restrict ourselves to those types actually needed for this particular publication.
    • MPolyPowersOfElement: powers of a collection of elements $f_{1},\dots,f_{k}\in P$;
    • MPolyComplementOfPrimeIdeal: complement of a prime ideal $Q\subset P$.
  • MPolyQuoLocRing: localizations $U^{-1}R$ of quotients $R=P/I$, considered as a finite $P$-algebra, at multiplicative sets $U\subset P$ in polynomial rings as above.
Polynomial rings and their quotients have already been implemented in OSCAR and its subsystems. Based on this work, the second named author developed the localizations.
Note that we work exclusively with multiplicative sets in polynomial rings $P$. In particular, localizations of quotient rings $R=P/I$ are realized as a localization of $R$, where we see $R$ as a finite $P$-module with a multiplication map. Successive localizations are automatically flattened to a single localization at products of multiplicative sets. There are various technical reasons supporting this decision; one of them being that quotient rings are in general not factorial while fast factorization of elements turns out to be key for performance of this implementation. From a purely mathematical viewpoint, these choices for the construction of localizations do not impose any relevant restrictions.
A common feature shared by all these four types of rings is that they posses coordinate functions, i.e. elements $x_{1},\dots,x_{n}\in R$ such that every element $f\in R$ is a fraction $f=\frac{p(x)}{q(x)}$ of polynomials $p,q\in\mathbf{k}[X_{1},\dots,X_{n}]$ evaluated at the $x_{i}$. Such coordinate functions are important as their images uniquely determine any homomorphism
$\varphi\colon R\to Q,\quad\varphi(x_{i})=g_{i}\in Q$
to an arbitrary commutative $\mathbf{k}$-algebra $Q$. By algebraic prolongation, an element $f=p/q$ as above is then taken to $\varphi(f)=p(g)\cdot q(g)^{-1}\in Q$. Based on this fact, homomorphisms from and to either of these four rings are supported in OSCAR, together with their common operations such as computing kernels or preimages of ideals.
6.2. Functionality for ideals and modules
For each one of the four types of rings, as in Section 6.1, we support the common functionality for ideals such as sums, intersections, radicals, minimal associated primes, primary decomposition, ideal quotients, saturation, and ideal membership, and similar functionality for finitely generated free modules and subquotients.
The generic OSCAR framework for finitely generated (graded) modules represented as subquotients, and its instance for modules over polynomial rings, is due to J. Böhm and his student A. Dinges, see also [Dinges]. The computational backend is Singular [Sin]. For the localizations and quotient rings the second named author has partially picked up on the ideas communicated in [Posur18] for so-called computable rings. Recall that a ring $R$ is computable if there are algorithms which solve the following three problems:
  1. (1)
    Given a morphism of free $R$-modules $\varphi\colon R^{m}\to R^{n}$, compute a generating set $v_{1},\dots,v_{r}\in R^{m}$ for the kernel of $\varphi$.
  2. (2)
    Given a finitely generated submodule $I=\langle v_{1},\dots,v_{r}\rangle\subset R^{m}$ and an element $w\in R^{m}$, decide whether $w\in I$.
  3. (3)
    Given an element $w\in I\subset R^{m}$ as above, find a set of elements $\lambda_{1},\dots,\lambda_{r}\in R$ such that $w=\sum_{i=1}^{r}\lambda_{i}\cdot v_{i}$.
Now most algorithms involving ideals and modules in Commutative Algebra can be reduced to these three building blocks. There is a layer of generic code for finitely presented modules in OSCAR doing that.
The key result in [Posur18] is that for a computable ring $R$, the localization at a multiplicative set $U\subset R$ is again computable if there is an algorithm for the following problem:
  1. (4)
    Given a finitely generated ideal $I\subset R$, decide whether $U\cap I$ is non-empty and if so, produce a single element $x\in U\cap I$.
Building on this result, the second named author has extended the generic code base for finitely presented modules and ideals so that it is now possible to have full functionality over localizations in OSCAR by only implementing one single method for a given new type of multiplicative sets.
However, there are several problems for which a direct reduction to the above three methods does not yield a reasonable performance, or, even worse, there is no known reduction to the above three methods. For these cases the julia programming language allows us to clip in custom implementations based on the multiple dispatch system. One example for a functionality which is not easily reduced to (1)-(3) above is primary decomposition which is a backbone for modeling of algebraic cycles and intersection theory. Another example where our actual implementations do not rely on (4) is localization at powers of elements $U=\{f^{k}:k\in\mathbb{N}_{0}\}$, $f\in R$. For these, working with saturations of ideals and modules has proven to be more effective.
6.3. Affine schemes
In general, an affine scheme in OSCAR is merely a wrapper for one of the above types of rings and a morphism of affine schemes is an arrow-reversing wrapper of the ring homomorphisms for the pullback of regular functions. The abstract type for affine schemes and their morphisms for which their minimal interfaces are defined, are AbsAffineScheme and AbsAffineSchemeMor, respectively. Minimal concrete implementations of these interfaces are realized for the concrete types AffineScheme and AffineSchemeMor. An ordinary affine scheme $X$ is constructed calling the function spec(R) on a ring $R$ of the above four types. The constructor for a morphism $X\to Y$ is morphism(X, Y, [g_1, ..., g_n]) where $g_{1},\dots,g_{n}\in R$ are the pullbacks of the coordinate functions for $Y$ on $X=\mathrm{Spec}\,R$.
6.3.1. Comparison of affine schemes
As already explained above, all of the four types of rings $R$ supported for affine algebraic schemes have a governing polynomial ring $P=\mathbf{k}[X_{1},\dots,X_{n}]$ in the background, the ambient_coordinate_ring. Note that these arise by the nature of their implementation; mathematically these polynomial rings depend on a choice of the coordinate functions. Even though this deviates from the strict mathematical definitions, we found it reasonable and practical to allow for comparison of either two affine schemes $X=\mathrm{Spec}\,R$ and $Y=\mathrm{Spec}\,R^{\prime}$ based on their embeddings $\mathrm{Spec}\,R\hookrightarrow\mathrm{Spec}\,P$ and $\mathrm{Spec}\,R^{\prime}\hookrightarrow\mathrm{Spec}\,P$ in their “ambient spaces” whenever both $R$ and $R^{\prime}$ are constructed using the same polynomial ring $P$ as an object in memory.
6.3.2. Inheritance within the schemes framework
Even though the julia programming language has no natural patterns for inheritance for its types, there are various possibilities to mimic type-inheritance. For our schemes framework we decided to implement the following simple pattern.
  • Introduce an abstract type MyAbstractType and declare an interface for that; including all getters my_getter(::MyAbstractType), but probably also more advanced functionality.
  • Implement the generic getters via
      my_getter(obj::MyAbstractType) = my_getter(underlying_concrete_instance(obj))
    and similarly for the other generic methods of the interface’s functions.
  • Introduce at least one minimal concrete type MyMinimalConcreteType <: MyAbstractType together with a full implementation of the interface by methods for this concrete type; i.e.
      my_getter(::MyMinimalConcreteType).
  • For further, more specialized concrete types NewConcreteType <: MyAbstractType, either implement the interface for MyAbstractType directly by overwriting the respective methods, or indirectly, by storing an instance of MyMinimalConcreteType inside NewConcreteType and forwarding of its functionality via implementation of underlying_concrete_instance(::NewConcreteType).
For instance, this allows us to introduce specialized types for closed embeddings as
  ClosedEmbedding <: AbsAffineSchemeMor.
Besides the usual data required for a morphism of affine schemes, this also holds the information on the ideal defining the image in its ambient scheme. While the multiple dispatch system of the julia language allows for more fine-grained handling of the functionality for such ClosedEmbeddings, forwarding a concrete instance of AffineSchemeMor in the internals enables the use of generic code written for AbsAffineSchemeMor and thus avoids code duplication.
Having described this pattern, and its importance for our implementation, we will in the following only briefly comment on the abstract types for which an interface is defined, and the concrete types for their minimal implementations. For further details, we refer the reader to the source code.
6.4. Covered schemes
Recall that a scheme over a ring $\mathbf{k}$ is a locally ringed space $(X,\mathcal{O}_{X})$ which admits an open cover $(U_{i})_{i\in I}$ such that each $(U_{i},\mathcal{O}_{X}|U_{i})\cong\mathrm{Spec}\,R_{i}$ as locally ringed spaces for some finitely generated $\mathbf{k}$-algebra $R_{i}$.
In OSCAR, we start from finitely many affine patches $(U_{i},\mathcal{O}_{X}|U_{i})_{i\in I}=(\mathrm{Spec}\,R_{i})_{i\in I}$ and isomorphisms $f_{i,j}$, called glue maps, between principal open subsets [^7] We have partial support for gluings along more general open subsets, but this is not required within the scope of this paper.:
[…]
where we require for the glue maps that $f_{j,k}\circ f_{i,j}=f_{i,k}$ on the overlap $U_{i,j}\cap f_{i,j}^{-1}(U_{j,k})$. On the mathematcal side, this means that we define
(18) $Y:=\bigsqcup_{i}U_{i},\qquad X:=Y/\sim$
where for $x\in U_{i}$ and $y\in U_{j}$ we set $x\sim y$ if $f_{i,j}(x)=y$. The structure sheaf $\mathcal{O}_{X}$ is the unique sheaf of rings such that the canonical projection $\pi\colon Y\to X$ is a morphism of locally ringed spaces.
Note that there is an implicit compromise in this construction: There is, for instance, no single version of $U_{i}\cap U_{j}$ for two affine charts $U_{i}$ and $U_{j}$ of $X$. Instead, we are practically dealing with charts on $Y$. In particular, $U_{i,j}$ and $U_{j,i}$ are not equal in memory (i.e. on $Y$) but both represent $U_{i}\cap U_{j}$. However, introducing a new type of affine scheme which would allow for different such representatives seemed excessive and would lead to inconsistencies with the design decisions made for affine schemes. Therefore we treat $U_{i,j}$ and $U_{j,i}$ in the above sense as different schemes a priori, but we allow to ask for the identifying maps $f_{ij}\colon U_{i,j}\to U_{j,i}$. This has further consequences down the road as we shall see in the next section on sheaves.
In general, it is the user’s responsibility to provide charts and gluings such that the resulting scheme is well defined. For most practical applications, however, the construction of a scheme will ensue from standard building blocks such as projective spaces or, more generally, toric varieties, and subschemes thereof, as well as projective bundles, blow-ups, etc. In all such cases the gluings are provided automatically and separatedness of the scheme is guaranteed throughout the process.
Remark 6.1.
The actual computation of automatically generated gluings can be rather expensive and, given that the total number of gluings of a scheme $X$ is quadratic in the number $N$ of affine charts, should in general be avoided if possible. Therefore, most gluings are implemented lazy, i.e. they are computed only on request and then cached.
The interface for gluings as in Equation 17 is declared for the abstract type AbsGluing and comprises patches, gluing_domains, and gluing_morphisms for the identifying maps in the middle. The concrete type of any covered scheme in OSCAR is a subtype of the abstract type AbsCoveredScheme and has at least one Covering, its default_covering. The latter then provides the affine patches and their pairwise gluings. A basic implementation of the interface specified for AbsCoveredScheme is realized for the concrete type CoveredScheme. Again, such basic implementations can be enriched with arbitrary additional data for specialized types and we have done so with our implementation of EllipticSurface. Among other things, the latter also provides its weierstrass_model, the associated elliptic curve over a function field via generic_fiber, the trivial_lattice, and the weierstrass_contraction to project to the Weierstrass model.
6.5. Coherent sheaves
Recall that a presheaf $\mathcal{F}$ on a topological space $X$ is a contravariant functor which associates to every open subset $U\subset X$ an object $\mathcal{F}(U)$ and to every open inclusion of open sets $U\hookrightarrow V$ a restriction morphism
$\rho_{V,U}\colon\mathcal{F}(V)\to\mathcal{F}(U).$
A presheaf $\mathcal{F}$ is a sheaf, if for any open $U\subset X$, any covering $\{U_{i}\}_{i\in I}$ of $U$, and any collection of sections $s_{i}\in\mathcal{F}(U_{i})$ with $s_{i}=s_{j}$ on $U_{i}\cap U_{j}$ there exists a unique global section $s\in\mathcal{F}(U)$ with $s_{i}=\rho_{U,U_{i}}(s)$.
A coherent sheaf on a covered scheme $X$ as above is a sheaf of $\mathcal{O}_{X}$-modules which are locally finitely generated. On an affine scheme $U=\mathrm{Spec}\,R$ any coherent sheaf $\mathcal{F}$ is uniquely determined by the finitely generated $R$-module $M=\mathcal{F}(U)$.
While the definition of these mathematical notions is quite straightforward, we are facing various constraints and challengens when attempting to model this in a computer:
  • We have to model a sheaf by a finite amount of data.
  • We do not have direct access to the topology of $X$, only to $Y$ as in (18) and the glue maps $f_{i,j}$.
In order to surmount these challenges, we model a coherent sheaf $\mathcal{F}$ on $X$ by its pullback along the canonical projection $Y\to X$. This forces us to make several compromises which lead to a practically working substitute for the notion of a coherent sheaf, but at the expense that we slightly deviate from the strict mathematical concept.
The first and probably most important compromise is that we actually only model presheaves, i.e. we provide the functions $U\mapsto\mathcal{F}(U)$ on admissible open subsets $U$ and $U\supset V\mapsto\rho_{U,V}$ for admissible open inclusions. For the concrete implementations, we leave it as the programmer’s responsibility to assure that they behave like honest sheaves.
Second, we restrict the form of open subsets $U\subset X$ which are recognized as admissible for the production of the module $\mathcal{F}(U)$ for a coherent sheaf $\mathcal{F}$; this will be detailed below.
The third compromise is inherent in our construction: the implicit identification of $U_{i}\subseteq Y$ with $\pi(U_{i})\subseteq X$. In consequence, objects which are mathematically the same on $X$, appear from different representations on $Y$ by concrete data in memory: For $U_{i},U_{j}$ open in $X$ the module $\mathcal{F}(U_{i}\cap U_{j})$ can be represented as both, the localization $\mathcal{F}(U_{i,j})$ of the module $\mathcal{F}(U_{i})$ and as localization $\mathcal{F}(U_{j,i})$ of $\mathcal{F}(U_{j})$. Both representations are reconciled by giving an isomorphism $\rho_{ij}\colon\mathcal{F}(U_{i,j})\to\mathcal{F}(U_{j,i})$ as part of the data. In particular this holds for the structure sheaf $\mathcal{O}_{X}$.
The admissible open subsets for a coherent sheaf $\mathcal{F}$ on a covered scheme $X$ are
  • the affine_charts of $X$, i.e. the patches $U_{i}$ of its default_covering $(U_{i})_{i\in I}$,
and by recursive extension
  • any PrincipalOpenSubset of an admissible open subset and
  • any SimplifiedAffineScheme of an admissible open subset.
The latter two types are special concrete subtypes of AbsAffineScheme which relate to an ambient_scheme, resp. to an original_scheme. A PrincipalOpenSubset arises as the complement of the zero locus of a single element $h\in\mathcal{O}_{X}(V)$, the complement_equation, on an affine scheme $V$.
Remark 6.2.
The SimplifiedAffineScheme was introduced in order to eliminate superfluous variables whenever possible. To this end, we apply Singular’s elimpart heuristic to the modulus $I$ of a quotient ring $R=P/I$ for $V=\mathrm{Spec}\,R$ and wrap the result in an instance of SimplifiedAffineScheme, together with the identifying maps between both representations of $\mathrm{Spec}\,R$. Since the complexity of Buchberger’s algorithm is doubly exponential in the number of variables, this is a crucial step in keeping things computable; especially in setups like iterated blowup constructions.
With this recursively defined structure of a “parent node” given by the ambient, resp. original scheme, the admissible open subsets of a covered scheme $X$ form a forest (i.e. a union of trees in the graph-theoretic sense), whose roots, the affine charts of $X$, are pairwise related by gluings. This structure can then be used to indicate how to actually obtain the module $\mathcal{F}(U)$ for admissible open subsets $U\subset X$ different from the patches of the default covering: If $U$ is a PrincipalOpenSubset, then $\mathcal{F}(V)$ is constructed as a localization of $\mathcal{F}(V)$ where $V$ is the ambient_scheme of $U$. Similarly, in case $U$ is a SimplifiedAffineScheme, the module $\mathcal{F}(U)$ is constructed via change of base rings from $\mathcal{F}(V)$ along the identifying morphisms with the original_scheme $V$. This pattern eventually recurses up until some patch is reached, on which a module for $\mathcal{F}$ is already known; be it by the original construction of $\mathcal{F}$, or a previous call.
Similarly, the tree structure is used to construct the restriction maps $\mathcal{F}(V)\to\mathcal{F}(U)$ for $U\subset V$ in $X$. If $U$ and $V$ come to lie in different affine charts, this involves the gluings on the root nodes of the respective two trees.
Remark 6.3.
The complement $V$ of $\{xy-z^{2}=0\}\subset\mathbb{P}^{2}$ is affine, but not contained in any of the standard affine charts of $\mathbb{P}^{2}$, seen as a covered scheme. To cover such cases, further types of admissible subsets may be added to OSCAR in the future.
The abstract type for which the user-facing interface for coherent sheaves is declared, is AbsCoherentSheaf. The overall design of coherent sheaves in OSCAR is (and has to be) lazy and cached: No instance $\mathcal{F}(U)$ is computed, unless required for the definition of a coherent sheaf or requested by an explicit call. In general, the programmer is free to choose how the modules $\mathcal{F}(U)$ and their restriction morphisms are produced for their concrete subtypes of AbsCoherentSheaf. While the constructor of the basic implementation requires given modules and identification maps over the affine charts of $X$ and their gluings, there are also special types like e.g. PullbackIdealSheaf, the constructor of which requires only an AbsCoveredSchemeMorphism and an AbsIdealSheaf on its codomain. Yet another type of ideal sheaf is PrimeIdealSheafFromChart whose constructor only needs one prime ideal $I\subset\mathcal{O}_{X}(U)$ on one non-empty admissible affine open $U\subset X$ as input and then extends to other charts from there. All of these types come with different patterns for their computation on admissible open subsets, realized by different sets of methods of the internal function produce_object. See the source code for details and more, concrete examples.
6.6. Morphisms of covered schemes
Let $X=\bigcup_{i\in I}U_{i}$ and $Y=\bigcup_{j\in J}V_{j}$ be covered schemes. In general, a morphism $f\colon X\to Y$ is given by means of a refinement $X=\bigcup_{i\in I,j\in J_{i}}U_{i,j}$ of the covering of $X$ and morphisms of affine schemes
(19) $f_{i,j}\colon U_{i,j}\to V_{j},$
which glue naturally on the respective overlaps. In OSCAR this information is held in a CoveringMorphism which is in turn wrapped in some concrete instance of AbsCoveredSchemeMorphism.
For many setups this already provides convenient data structures to work with. For instance for the blow-up $\pi\colon\mathrm{Bl}_{\mathcal{I}}X\to X$ of an ideal sheaf $\mathcal{I}$ on a covered scheme $X$ the standard constructions of $\mathrm{Bl}_{\mathcal{I}}$ as a covered scheme from $X=\bigcup_{i\in I}U_{i}$ leads to a covering $\mathrm{Bl}_{\mathcal{I}}=\bigcup_{j\in J}V_{j}$ of $Y$ for which one does not need a refinement to write down the local morphisms $V_{j}\to U_{i(j)}$ of affine schemes for the projection map.
This example occurs in our implementation of elliptic surfaces when we construct a relatively minimal model $X$ as the minimal resolution of its (minimal) Weierstrass model $W$. The resolution $X\to W$ is a sequence of blowups and, in particular, an AbsCoveredSchemeMorphism.
6.7. Rational functions
A variety is a geometrically integral, separated scheme of finite type over a field $\mathbf{k}$. The corresponding type in OSCAR is AbsCoveredVariety. To a variety $X$ we can associate its field of rational functions
$k(X)=\mathrm{Quot}(\mathcal{O}_{X}(U))$
which is isomorphic to the fraction field of the ring $\mathcal{O}_{X}(U)$ of regular functions of any nonempty affine open subset $U\subset X$. Consequently, an element $f\in k(X)$ comes with many representatives: At least one for every affine chart of $X$.
The type for elements of $k(X)$ is VarietyFunctionFieldElem. Any such element can be asked to produce a representative on a given affine open subset $V\subset X$ which is admissible in the sense explained in the previous section on coherent sheaves. Representatives are provided as fractions of elements in the ambient_coordinate_ring; a decision owed to the fact that we do not support fraction fields of integral quotients of polynomial rings in OSCAR.
The implementation of rational functions on $X$ is lazy in the sense that representatives are computed only in the moment where the user asks for one. At first, only one representative $(U,f)$ with $f=\frac{p}{q}$ for one affine chart $U\subset X$ is stored in the data structure for $f$. Once the user asks for a representative on an admissible open subset $V\subset X$, a heuristic assessement of the presumed complexity of the production from already existing representatives $(U^{\prime},f^{\prime})$ is made. Such production involves evaluation of the gluing of $U^{\prime}$ with $V$ which must be assumed to be computationally expensive in general.
6.8. Birational Maps
For morphisms like the sought for automorphism of an elliptically fibered K3 surface $f_{OY}\colon X_{OY}\to X_{OY}$ it is prohibitively expensive to compute a covering of $X_{OY}$ which is sufficiently fine-grained so that $f_{OY}$ can be described by local maps on its patches. However, this is not needed for our purposes such as computing the pushforward of divisors. It is therefore better to view $f_{OY}$ as a birational map. Let us recall the definition:
Let $X,Y$ be varieties, $U,U^{\prime}\subseteq X$ and $V,V^{\prime}\subseteq Y$ be non-empty open subsets. Two morphisms $\varphi\colon U\to V$ and $\varphi^{\prime}\colon U^{\prime}\to V^{\prime}$ are said to be equivalent if they agree on $U\cap U^{\prime}$ (which is non-empty because $X$ is a variety). The corresponding equivalence class is called a birational map and it is called dominant if the image of any representative is dense in $Y$.
Within the scope of this paper, we work with dominant birational maps $f\colon X\dashrightarrow Y$ which extend (uniquely) to a morphism $X\to Y$, that is, which have a trivial indeterminacy locus. We nevertheless use birational morphisms and their representatives in the internals of the implementation, as we shall see below.
A dominant birational map $f\colon X\dashrightarrow Y$ is uniquely determined by its pullback of function fields $k(Y)\to k(X)$ and this is also the input data provided from our computations: Given the Weierstrass chart $W_{Y}$ on an elliptic K3 surface $Y$ with coordinates $(t,x,y)$, the computations from the $2$-neighbor step provide us with a Weierstrass chart $W_{X}$ of $X$ and the pullbacks $f^{*}(t),f^{*}(x)$, and $f^{*}(y)$ as rational functions on $W_{X}$. Since $K_{X}$ is nef and $X$ a (smooth) surface, we know a priori that $f\colon X\to Y$ is an isomorphism (although it is represented as a rational map).
All this is incorporated in a specialized type for AbsCoveredSchemeMorphism called MorphismFromRationalFunctions. An approach similar to ours, but with a view towards applications in massively parallel setups, has also been pursued and implemented by B. Mirgain in Singular; see [Mirgain24]. Any concrete instance of this type does not attempt to compute a representation $\{f_{i,j}\colon U_{ij}\to V_{j}\}$ of itself as in Equation 19, unless requested by the user. The functionality we need for such an automorphism –basically pullback and pushforward of divisors– is mostly overwritten by methods which do not require the computation of a full CoveringMorphism. This is crucial to the success of our computations, as realizing $f_{OY}$ via a full-fledged CoveringMorphism will probably not terminate in practice.
6.9. Specialization
Recall that the base change of $X/R$ along a morphism $R\to k$, is defined by the fiber product
${X\times\mathrm{Spec}\,k}$${X}$${\mathrm{Spec}\,(k)}$${\mathrm{Spec}\,(R).}$$\scriptstyle{\mathrm{red}}$
OSCAR can compute the fiber product and the corresponding morphisms of AbsCoveredSchemes by looking at the individual charts and their gluings. By functoriality of the fiber product, local constructions patch together to give well-defined global objects.
In our application, we are facing the following caveat. The instances of AbsCoveredSchemes that we are working with are EllipticSurfaces which hold significantly more information than just a covered scheme, and in positive characteristic we would like to have the counterparts of this information available again. In particular, recall from Section 4.4 that we need access to the specialization map $\mathrm{sp}\colon\mathrm{NS}(X)\to\mathrm{NS}(X_{k})$, so a “forgetful” reduction map which only has a CoveredScheme in its codomain will not suffice.
Now we could attempt to overwrite the method for reduction to positive characteristic for an EllipticSurface, but that would be technically very challenging, as we would have to dynamically adjust all cached objects such as, for instance, the Weierstrass models and projections, the basis of the algebraic_lattice, etc. Instead, we pursue a different approach. Given an EllipticSurface $X$ over a ring $R$ of characteristic zero and a reduction map $R\to k$, we first specialize the Weierstrass chart $W$ of $X$ to obtain the Weierstrass chart $W_{k}$ of $X_{k}$ and construct the latter as an a-priori independent EllipticSurface. Via the identification
$X\times\mathrm{Spec}\,k\hookleftarrow W_{k}\hookrightarrow X_{k}$
we obtain a birational map
$X\times\mathrm{Spec}\,k\dashrightarrow X_{k}$
(which is in fact an isomorphism) defined as a MorphismFromRationalFunctions. Then we obtain the specialization map as the pushforward of the composition
$X\overset{\mathrm{red}}{\longleftarrow}X\times\mathrm{Spec}\,k\overset{\cong}{\longrightarrow}X_{k}$
On the true EllipticSurface $X_{k}$ we can now identify any given divisor as an element in $\mathrm{NS}(X_{k})$ again and we get an identification of the generators for the algebraic lattices on both $X$ and $X_{k}$. Similarly, we obtain $(f_{OY})_{k}$ as a MorphismFromRationalFunctions by specialization of its representation as a rational map $W\dashrightarrow W$ of the Weierstrass chart only.
Another advantage of this approach over the direct one is that data for $X_{k}$, such as the gluings between charts of $X_{k}$, can be computed directly in positive characteristic rather than in characteristic 00 (which is slower) and performing a base change.
6.10. Algebraic cycles and divisors
Following [Fulton98], an algebraic cycle of dimension $k$ on a scheme $X$ is a formal linear combination [^8] Unless otherwise specified, we will assume integer coefficients throughout.
$C=\sum_{i=1}^{N}c_{i}\cdot W_{i}$
of irreducible subvarieties $W_{i}\subset X$ of dimension $k$. While in mathematical theory it comes at no expense to make such irreducibility assumptions on the summands of an algebraic cycle, this is not practical in an implementation; mostly due to the computational cost of primary decomposition. Moreover, while it is theoretically possible to introduce a new concrete type of AbsCoveredScheme for subschemes or -varieties of some ambient scheme, we have not taken this step, yet. Instead, we decide to model an algebraic cycle as a formal linear combination of ideal sheaves
$C=\sum_{i=1}^{N}c_{i}\cdot\mathcal{I}_{i}.$
with $\mathcal{I}_{i}$ representing a subscheme $Y_{i}$ of $X$ which is not necessarily reduced or irreducible. If needed, the user can transform any such cycle following [Fulton98, Section 1.5] via irreducible_decomposition into a linear combination
$C=\sum_{\mathcal{P}\textnormal{ prime}}\left(\sum_{i}c_{i}\cdot\operatorname{length}_{\mathcal{O}_{X,\mathcal{P}}}(\mathcal{O}_{X,\mathcal{P}}/(\mathcal{I}_{i})_{\mathcal{P}})\right)\cdot\mathcal{P}$
where the sum ranges over all prime ideal sheaves $\mathcal{P}$, i.e. ideal sheaves whose scheme theoretic zero loci are irreducible subvarieties of $X$ and $\mathcal{O}_{X,\mathcal{P}}$ is the local ring at $\mathcal{P}$ (isomorphic to $(\mathcal{O}_{X}(U))_{\mathcal{P}(U)}$ for any affine open $U\subset X$ is any affine chart with $\mathcal{P}(U)\neq\langle 1\rangle$).
The user interface for algebraic cycles on covered schemes in Oscar is declared for the abstract type AbsAlgebraicCycle. The basic implementation is based on a dictionary, the keys of which are the components $\mathcal{I}_{i}$, pointing to their coefficients $c_{i}$. If $X$ itself is equidimensional, a Weil divisor is an algebraic cycle of codimension one. They basically share their functionality with the algebraic cycles but they come with their own abstract type AbsWeilDivisor.
For Cartier divisors we slightly deviate from the common description in the literature and implement a CartierDivisor as a formal sum of EffectiveCartierDivisors; see e.g. the Stacks project. An EffectiveCartierDivisor is internally based on an ideal sheaf $\mathcal{I}$ which is locally principal for some non-zerodivisor. In particular, any EffectiveCartierDivisor can be asked for the trivializing_covering and for every patch $V$ of that covering the ideal $\mathcal{I}(V)$ is then of the desired form.
Within the scope of this paper Cartier divisors are naturally encountered as exceptional divisors from the resolution of singularities of the Weierstrass model of an elliptically fibered K3 surface. Every Cartier divisor can be transformed into a Weil divisor, but in general not the other way around.
6.11. Pushforward and pullback of divisors
Weil divisors or, more generally, algebraic cycles are covariant objects for proper maps while Cartier divisors correspond – up to rational equivalence – to algebraic line bundles and are therefore of contravariant nature. In OSCAR we have partial support for the associated pushforward and pullback functionality for AbsCoveredSchemeMorphisms. The general pushforward of algebraic cycles for proper maps requires the computation of the degree of a field extension of fields of rational functions on varieties (cf. [Fulton98, Section 1.4]), a functionality which is not yet available in OSCAR.
In this paper we are mostly dealing with isomorphisms of varieties which are modeled as instances of MorphismFromRationalFunctions. In this setting pushforward and even pullback of algebraic cycles is straightforward and can be reduced to mapping (non-closed) points of $X$.
The implementation of the pushforward of an algebraic cycle $D$ along a MorphismFromRationalFunctions $f\colon X\to Y$ is based on the following observation. Let $\mathcal{P}$ be a prime ideal sheaf on $X$. If $U\subset X$ is any affine chart of $X$ with $\mathcal{P}(U)\neq\langle 1\rangle$, then the open subsets $U\cap f^{-1}(V_{j})$ cover $U$ as $V_{j}$ ranges through the affine charts of $Y$. Hence, there must be at least one $V=V_{j}$ for which $\mathcal{P}(U\cap f^{-1}(V))\neq\langle 1\rangle$. In case $U\cap f^{-1}(V)$ is not principally open in $U$, we may restrict to some principally open subset $W\subset U\cap f^{-1}(V)$ with $\mathcal{P}(W)\neq\langle 1\rangle$. Once we found such a pair $(U,V)$, we can realize a local representative $f\colon U\cap f^{-1}(V)\to V$ of $f$, compute the scheme-theoretic image of the point $\mathcal{P}$ along this morphism of affine schemes and return the corresponding ideal sheaf on $Y$.
Given that the rational map $f$ in our application is huge, the practical difficulty is to find a suitable pair $(U,V)$ such that the computation of the image $f(\mathcal{P}(U))$ terminates. Thus, we proceed as follows:
  1. (1)
    Assume that the divisor is of the form $D=\sum_{i}d_{i}\cdot\mathcal{P}_{i}$ for prime ideal sheaves $\mathcal{P}_{i}$ and reduce by linearity to the pushforward of a prime component $\mathcal{P}=\mathcal{P}_{i}$.
  2. (2)
    Find one affine chart $U\subseteq X$ in the default_covering of $X$ where $\mathcal{P}$ is visible, i.e. where $\mathcal{P}(U)\neq\langle 1\rangle$.
  3. (3)
    Iterate through the patches $V_{j}\subseteq Y$ of the default_covering of the codomain $Y$ of $f$ and compute the pullbacks of the coordinate functions $y_{1},\dots,y_{k}$ for the ring $\mathcal{O}_{Y}(V_{j})$ as rational functions on $U$ for all $V_{j}$.
  4. (4)
    Sort the codomain patches $V_{j}$ based on the complexity of the rational functions for $f$ on the pairs $(U,V_{j})$ and iterate through the $V_{j}$ in that order. Realize $f$ on some principal open subset $W=D(h)\subset U\cap f^{-1}(V)$ as a morphism of affine schemes
    $f\colon W\to V.$
    The element $h\in\mathcal{O}_{X}(U)$ arises from the denominators of the fractions for the local representation of $f$ on this pair of patches $(U,V)$. If $\mathcal{P}(W)\neq\langle 1\rangle$, compute the scheme theoretic image of this point and return the associated prime ideal sheaf on $Y$.
  5. (5)
    If the previous step fails, iterate through the $V_{j}$ again. For every $V=V_{j}$ compute the maximal Zariski open subset $W^{\prime}\subset U$ which allows for a realization of $f\colon W^{\prime}\to V$ as a morphism of schemes. Such $W^{\prime}$ can not be expected to be principally open in $U$, but it can be covered by principal open subsets. Eventually we must hit some $W^{\prime}$ for which $\mathcal{P}(W^{\prime})\neq\langle 1\rangle$ and we can finish as above.
Correctness of this algorithm then follows from the realizability of $f$ as a morphism of covered schemes.
Pullback of Weil divisors along an isomorphism $f\colon X\to Y$ given as a MorphismFromRationalFunctions is implemented similarly. Depending on the use case, sometimes the pushforward is faster, sometimes the pullback.
For an automorphism $f$ of an EllipticSurface $X$ we have implemented an even more specialized method for the pushforward of the known algebraic_lattice of $X$. For this we make a pre-assessement to determine the nature of the image of any component. This is achieved by composing $f$ with the projection $\pi\colon X\to\mathbb{P}^{1}$ of the elliptic fibration and pushforward of the components to $\mathbb{P}^{1}$: If their dimension is one, the respective component is mapped to a (multi-) section in $X$, if the dimension is zero, it is taken to a fiber component. This information can be used to significantly narrow down the affine charts $V$ in the codomain of $f$ for which one has to create local realizations of $f$ and we get significant speedups compared to the generic method for pushforward.
6.12. Computation of intersection numbers
On an elliptically fibered K3 surface $X$ computations of intersection numbers are important for the Gram-matrix of the algebraic_lattice and for representing the numerical equivalence class of an arbitrary divisor $D$ as a linear combination of a chosen generating set of the Néron-Severi group of $X$. Let us first take a look in greater generality:
Following [Fulton98, Definition 2.4.1 and Section 2.5], the natural intersection product on a variety $X$ over $\mathbf{k}$ is a is defined for a Cartier divisor [^9] In fact, Fulton introduces a pseudo-divisor to address various technical difficulties; we shall not go into these details here. $C$ via the first Chern class of the associated line bundle, seen as a map
$c_{1}(C)\colon Z_{k}(X)\to A_{k-1}(X),\quad D\mapsto c_{1}(C)\cap D$
from the $k$-cycles on $X$ to the Chow group of $(k-1)$-cycles on $X$. The integral of a zero-dimensional cycle $D=\sum_{\mathcal{P}}c_{\mathcal{P}}\cdot\mathcal{P}\in Z_{0}(X)$ is
(20) $\int_{X}D=\sum_{\mathcal{P}}c_{\mathcal{P}}\cdot[k(P):\mathbf{k}]$
where we assume all $\mathcal{P}$ to be prime. For an (equidimensional) surface $X$ which is proper over $\mathbf{k}$ one can then define the intersection number of a Cartier and a Weil divisor $C$ and $D$ as
$C.D=\int_{X}c_{1}(C)\cap D.$
Whenever $X$ is smooth, every Weil divisor is actually Cartier, so that the definition extends to pairs of Weil divisors. In this case, it is customary to declare the intersection of two Weil divisors to be this number, provided they are in sufficiently general position relative to each other; cf. [Hartshorne77, Chapter V, Theorem 1.1]. Since intersection numbers do not change under linear equivalence, this condition can always be achieved by a moving lemma, in theory at least.
While the theory is straightforward and well developed, one quickly reveals severe limitations of the above definitions when it comes to their practical feasibility in applications like ours. One of the main bottlenecks is again the requirement to have a Weil divisor as a linear combination of irreducible components. Fortunately, the computation of such a representation of a given cycle of dimension zero can be avoided if one is only interested in the integral of that cycle. We will describe this in more detail below.
Another instance where irreducibility of components is relevant in intersection theory is self-intersection of Cartier divisors, or, more generally, dealing with pairs of a Cartier and a Weyl divisor $C$ and $D$ where some irreducible component of $D$ happens to be properly contained in the support $|C|$ of the Cartier divisor $C$. Usually one employs some kind of “moving lemma” to resolve such situations. While we do have the infrastructure necessary to implement such moving lemmas, we shall not need them for the purpose of this article: On our elliptically fibered K3 surfaces all relevant divisors in the algebraic lattice are already given as irreducible curves and their self intersection is completely determined by their genus via the adjunction formula and the Riemann-Roch theorem for curves. Thus we deliberately hardcode such excess- and self-intersections whenever encountered in our specialized methods for an EllipticSurface.
In order to bring the definition of the integral, Equation (20), in a workable form for computation of intersection numbers on surfaces, we start with the following observations. Fix a component $\mathcal{P}$ of $D$ and let $P=V(\mathcal{P})\subset X$ be its zero locus; choose an affine chart $U\subset X$ for which $\mathcal{P}(U)\neq\langle 1\rangle$. Then the degree of the field extension
$[k(P):\mathbf{k}]=\dim_{\mathbf{k}}\mathcal{O}_{X}(U)/\mathcal{P}(U)$
is nothing but a vector space dimension of a quotient ring. Now if $\mathcal{Q}$ is $\mathcal{P}$-primary, then $[\mathcal{Q}]=n_{\mathcal{P}}\cdot[\mathcal{P}]$ with
$n_{\mathcal{P}}=\mathrm{length}_{\mathcal{O}_{X,P}}\mathcal{O}_{X,P}/\mathcal{Q}$
the length of the Artinian quotient ring $\mathcal{O}_{X,P}/\mathcal{Q}$; cf. [Fulton98, Chapter 1, Section 1.5]. Since we are working with zero dimensional cycles, it is easy to see that then
$n_{P}\cdot[k(P):\mathbf{k}]=\dim_{\mathbf{k}}\mathcal{O}_{X}(U)/\mathcal{Q}(U)$
so that we can compute the full contribution of $\mathcal{Q}$ to the integral directly without extracting the prime component $\mathcal{P}$.
For an arbitrary zero-dimensional ideal sheaf $\mathcal{I}$ on $X$ we observe that according to the Chinese Remainder Theorem the quotient $\mathcal{O}_{X}(U)/\mathcal{I}(U)$ decomposes as a direct sum
$\mathcal{O}_{X}(U)/\mathcal{I}(U)\cong\bigoplus_{i}\mathcal{O}_{X}(U)/Q_{i}$
for the primary components $Q_{i}$ of $\mathcal{I}(U)$ in every affine chart $U$ of $X$. Therefore the contribution to an integral $\int_{X}1\cdot\mathcal{I}$ from the components visible in a fixed affine chart $U\subset X$ is
$\dim_{\mathbf{k}}\mathcal{O}_{X}(U)/\mathcal{I}(U).$
This boils down to the computation of one single Gröbner basis rather than primary decompositions and computations of lengths.
At this point the main technical challenge is to avoid overcounting: One can compute contributions to the integral $\int_{X}D$ from every one of the affine charts $U$ of $X$, but in general the irreducible components $\mathcal{P}$ of $D$ will be visible in more than just one chart.
Without loss of generality we may assume that $D=1\cdot\mathcal{I}$ is given by a single ideal sheaf $\mathcal{I}$ on $X$. In principle, it is possible to compute local primary decompositions of $\mathcal{I}(U)$ and proceed with matching of the components via the gluing maps of either two charts. However, this turns out to be computationally rather involved with inferior performance: Not only is primary decomposition expensive, but also the computation of gluings for the matching of local components should be avoided.
To circumvent this problem, we write $X$ as a finite disjoint union of locally closed subsets $C_{i}$ which are contained as closed affine subschemes in the affine charts $U_{i}$ of $X$. In other words, we view
$X=\bigsqcup_{i}C_{i}$
as a constructible set. Note that there is no natural choice for such decompositions, but for most constructions it is possible to write them down. A common example for this is the decomposition
$\mathbb{P}^{n}=\mathbb{A}^{n}\sqcup\mathbb{P}^{n-1}=\mathbb{A}^{n}\sqcup\mathbb{A}^{n-1}\sqcup\cdots\sqcup\mathbb{A}^{1}\sqcup\mathbb{A}^{0}$
of projective space into the disjoint union of affine spaces, starting with the horizon $\mathbb{P}^{n-1}\subset\mathbb{P}^{n}$ at infinity and its complement $\mathbb{A}^{n}$ and the inductive continuation of this procedure with $\mathbb{P}^{n-1}$. Since the preimage of a constructible set under a morphism of schemes is constructible, we can pull back such a decomposition along sequences of blowups, forming total spaces of projective bundles, taking subschemes or finite covers and refine the decomposition further if necessary. This is done automatically in OSCAR.
Technically, this is modeled by what we colloquially call decomposition_info. For a given Covering $(U_{i})_{i\in I}$ of a covered scheme $X$ the decomposition info is a dictionary which holds for every patch $U_{i}$ of the covering a list of elements $f_{1,i},\dots,f_{k_{i},i}\in\mathcal{O}_{X}(U)$. These elements are chosen so that their common zero loci $C_{i}:=V(f_{1,i},\dots,f_{k_{i},i})\subset U_{i}\subset X$ provide a decomposition of $X$ as a disjoint union of locally closed subsets.
For the computation of an integral the decomposition_info comes in precisely to avoid overcounting: The contribution to $\int_{X}1\cdot\mathcal{I}$ from a single chart $U_{i}$ is computed as follows. We throw away all components of $I=\mathcal{I}(U)$ whose support does not lie in $C_{i}$. This only needs saturation and can be done without computing a primary decomposition. We obtain an ideal $J_{i}\subset\mathcal{O}_{X}(U_{i})$ for the remaining components and the contribution to the integral from this chart $U_{i}$ then amounts to $\dim_{\mathbf{k}}\mathcal{O}_{X}(U_{i})/J_{i}$.
7. Appendix
$\begin{matrix}\hline\cr\left(\begin{matrix}(\frac{1}{2}\cdot\alpha^{7}+\frac{5}{2}\cdot\alpha^{3}-\alpha^{2}-\alpha)\cdot t^{4}\\ +(\frac{1}{2}\cdot\alpha^{7}+\frac{1}{2}\cdot\alpha^{5}+\alpha^{4}+\frac{7}{2}\cdot\alpha^{3}-\alpha^{2}-\frac{1}{2}\cdot\alpha)\cdot t^{3}\\ +(-\frac{1}{2}\cdot\alpha^{6}+\alpha^{4}+2\cdot\alpha^{3}-\frac{1}{2}\cdot\alpha^{2})\cdot t^{2}\\ +(\frac{1}{2}\cdot\alpha^{7}-\frac{1}{2}\cdot\alpha^{6}-\frac{1}{2}\cdot\alpha^{5}-\frac{1}{2}\cdot\alpha^{4}+\frac{7}{2}\cdot\alpha^{3}-\frac{1}{2}\cdot\alpha^{2}+\frac{1}{2}\cdot\alpha-\frac{1}{2})\cdot t\\ +\frac{1}{2}\cdot\alpha^{7}-\frac{1}{2}\cdot\alpha^{4}+\frac{5}{2}\cdot\alpha^{3}+\alpha-\frac{1}{2}\end{matrix}\right.,&\left.\begin{matrix}(\frac{1}{2}\cdot\alpha^{6}+\frac{1}{2}\cdot\alpha^{5}+\frac{1}{2}\cdot\alpha^{4}+\frac{3}{2}\cdot\alpha^{2}-\frac{1}{2}\cdot\alpha-\frac{1}{2})\cdot t^{6}\\ +(-\alpha^{7}+\alpha^{5}+3\cdot\alpha^{4}-\alpha^{3}+3\cdot\alpha^{2}+\alpha)\cdot t^{5}\\ +(-\frac{3}{2}\cdot\alpha^{7}-2\cdot\alpha^{6}-3\cdot\alpha^{5}-\frac{3}{2}\cdot\alpha^{3}+4\cdot\alpha^{2})\cdot t^{4}\\ +(\alpha^{7}-\frac{1}{2}\cdot\alpha^{6}-5\cdot\alpha^{5}-\frac{15}{2}\cdot\alpha^{4}+\alpha^{3}+\frac{9}{2}\cdot\alpha^{2}-\alpha-\frac{1}{2})\cdot t^{3}\\ +(\frac{3}{2}\cdot\alpha^{7}+\frac{5}{2}\cdot\alpha^{6}-\frac{9}{2}\cdot\alpha^{4}-\frac{3}{2}\cdot\alpha^{3}+\frac{7}{2}\cdot\alpha^{2}+\frac{1}{2})\cdot t^{2}\\ +(\frac{3}{2}\cdot\alpha^{6}+2\cdot\alpha^{5}+\frac{3}{2}\cdot\alpha^{4}-2\cdot\alpha^{3}+\frac{3}{2}\cdot\alpha^{2}+\frac{3}{2})\cdot t\\ +\frac{1}{2}\cdot\alpha^{5}+\alpha^{4}+\alpha^{3}+\frac{1}{2}\cdot\alpha+1\end{matrix}\right)\\ \hline\cr\left(\begin{matrix}-t^{4}+(\frac{1}{2}\cdot\alpha^{6}+\frac{5}{2}\cdot\alpha^{2}-1)\cdot t^{2}\\ +\frac{1}{2}\cdot\alpha^{6}+\frac{5}{2}\cdot\alpha^{2}\end{matrix},\right.&\left.\begin{matrix}(-\frac{1}{2}\cdot\alpha^{7}-\frac{1}{2}\cdot\alpha^{5}-\frac{5}{2}\cdot\alpha^{3}-\frac{5}{2}\cdot\alpha)\cdot t^{5}\\ +(\frac{1}{2}\cdot\alpha^{7}-\frac{1}{2}\cdot\alpha^{5}+\frac{7}{2}\cdot\alpha^{3}-\frac{7}{2}\cdot\alpha)\cdot t^{3}+(\alpha^{7}+6\cdot\alpha^{3}-\alpha)\cdot t\end{matrix}\right)\\ \hline\cr\left(t^{4}-1,\right.&\left.\begin{matrix}(-\frac{1}{2}\cdot\alpha^{6}-\frac{7}{2}\cdot\alpha^{2})\cdot t^{4}+\frac{1}{2}\cdot\alpha^{6}+\frac{7}{2}\cdot\alpha^{2}\end{matrix}\right)\\ \hline\cr\left(\begin{matrix}(\alpha^{7}+\alpha^{6}+\frac{1}{2}\cdot\alpha^{5}+6\cdot\alpha^{3}+6\cdot\alpha^{2}+\frac{5}{2}\cdot\alpha)\cdot t^{4}\\ +(-\frac{7}{2}\cdot\alpha^{7}-\alpha^{6}+\frac{1}{2}\cdot\alpha^{5}+\alpha^{4}-\frac{41}{2}\cdot\alpha^{3}-6\cdot\alpha^{2}+\frac{7}{2}\cdot\alpha+6)\cdot t^{3}\\ +(-\frac{5}{2}\cdot\alpha^{6}-2\cdot\alpha^{5}-\alpha^{4}-\frac{29}{2}\cdot\alpha^{2}-12\cdot\alpha-6)\cdot t^{2}\\ +(\frac{7}{2}\cdot\alpha^{7}+\frac{5}{2}\cdot\alpha^{6}+\frac{1}{2}\cdot\alpha^{5}-\frac{1}{2}\cdot\alpha^{4}+\frac{41}{2}\cdot\alpha^{3}+\frac{29}{2}\cdot\alpha^{2}+\frac{7}{2}\cdot\alpha-\frac{5}{2})\cdot t\\ -\alpha^{7}+\frac{1}{2}\cdot\alpha^{5}+\frac{1}{2}\cdot\alpha^{4}-6\cdot\alpha^{3}+\frac{5}{2}\cdot\alpha+\frac{5}{2}\end{matrix},\right.&\left.\begin{matrix}(\frac{7}{2}\cdot\alpha^{7}+\frac{3}{2}\cdot\alpha^{6}-\frac{1}{2}\cdot\alpha^{4}+\frac{41}{2}\cdot\alpha^{3}+\frac{17}{2}\cdot\alpha^{2}-\frac{7}{2})\cdot t^{6}\\ +(-5\cdot\alpha^{7}+3\cdot\alpha^{6}+5\cdot\alpha^{5}+3\cdot\alpha^{4}-29\cdot\alpha^{3}+18\cdot\alpha^{2}+29\cdot\alpha+18)\cdot t^{5}\\ +(-18\cdot\alpha^{7}-16\cdot\alpha^{6}-\frac{15}{2}\cdot\alpha^{5}-105\cdot\alpha^{3}-94\cdot\alpha^{2}-\frac{87}{2}\cdot\alpha)\cdot t^{4}\\ +(29\cdot\alpha^{7}+\frac{15}{2}\cdot\alpha^{6}-5\cdot\alpha^{5}-\frac{15}{2}\cdot\alpha^{4}+169\cdot\alpha^{3}+\frac{89}{2}\cdot\alpha^{2}-29\cdot\alpha-\frac{89}{2})\cdot t^{3}\\ +(\frac{23}{2}\cdot\alpha^{6}+\frac{21}{2}\cdot\alpha^{5}+\frac{9}{2}\cdot\alpha^{4}+\frac{133}{2}\cdot\alpha^{2}+\frac{123}{2}\cdot\alpha+\frac{55}{2})\cdot t^{2}\\ +(-12\cdot\alpha^{7}-\frac{15}{2}\cdot\alpha^{6}-2\cdot\alpha^{5}+\frac{3}{2}\cdot\alpha^{4}-70\cdot\alpha^{3}-\frac{87}{2}\cdot\alpha^{2}-12\cdot\alpha+\frac{15}{2})\cdot t\\ +\frac{5}{2}\cdot\alpha^{7}-\alpha^{5}-\alpha^{4}+\frac{29}{2}\cdot\alpha^{3}-6\cdot\alpha-5\end{matrix}\right)\\ \hline\cr\left(\begin{matrix}(\frac{1}{2}\cdot\alpha^{4}+\frac{1}{2})\cdot t^{4}+(-\frac{1}{2}\cdot\alpha^{4}+\alpha^{2}-\frac{1}{2})\cdot t^{2}-\alpha^{2}\end{matrix},\right.&\left.\begin{matrix}(-\frac{1}{2}\cdot\alpha^{6}-\frac{3}{2}\cdot\alpha^{2})\cdot t^{6}+(\frac{1}{2}\cdot\alpha^{6}-\alpha^{4}+\frac{1}{2}\cdot\alpha^{2})\cdot t^{4}\\ +(\frac{3}{2}\cdot\alpha^{4}+\frac{1}{2})\cdot t^{2}-\frac{1}{2}\cdot\alpha^{4}+\alpha^{2}-\frac{1}{2}\end{matrix}\right)\\ \hline\cr\left(\begin{matrix}(\frac{1}{2}\cdot\alpha^{7}+\frac{5}{2}\cdot\alpha^{3}+\alpha^{2}-\alpha)\cdot t^{4}\\ +(\frac{1}{2}\cdot\alpha^{7}-\frac{1}{2}\cdot\alpha^{6}+\frac{1}{2}\cdot\alpha^{5}+\frac{1}{2}\cdot\alpha^{4}+\frac{3}{2}\cdot\alpha^{3}-\frac{1}{2}\cdot\alpha^{2}+\frac{3}{2}\cdot\alpha+\frac{1}{2})\cdot t^{3}\\ +(-\frac{1}{2}\cdot\alpha^{6}+\alpha^{4}-2\cdot\alpha^{3}-\frac{1}{2}\cdot\alpha^{2})\cdot t^{2}\\ +(\frac{1}{2}\cdot\alpha^{7}-\frac{1}{2}\cdot\alpha^{5}+\alpha^{4}+\frac{3}{2}\cdot\alpha^{3}+\alpha^{2}-\frac{3}{2}\cdot\alpha)\cdot t\\ +\frac{1}{2}\cdot\alpha^{7}+\frac{1}{2}\cdot\alpha^{4}+\frac{5}{2}\cdot\alpha^{3}+\alpha+\frac{1}{2}\end{matrix},\right.&\left.\begin{matrix}(-\frac{1}{2}\cdot\alpha^{7}+\frac{1}{2}\cdot\alpha^{6}-\frac{1}{2}\cdot\alpha^{4}-\frac{3}{2}\cdot\alpha^{3}+\frac{3}{2}\cdot\alpha^{2}+\frac{1}{2})\cdot t^{6}\\ +(-\alpha^{7}+\alpha^{5}-3\cdot\alpha^{4}-\alpha^{3}-3\cdot\alpha^{2}+\alpha)\cdot t^{5}\\ +(-\frac{3}{2}\cdot\alpha^{7}+3\cdot\alpha^{5}-7\cdot\alpha^{4}-\frac{3}{2}\cdot\alpha^{3}-3)\cdot t^{4}\\ +(-\alpha^{7}-\frac{1}{2}\cdot\alpha^{6}+5\cdot\alpha^{5}-\frac{15}{2}\cdot\alpha^{4}-\alpha^{3}+\frac{9}{2}\cdot\alpha^{2}+\alpha-\frac{1}{2})\cdot t^{3}\\ +(-\frac{5}{2}\cdot\alpha^{6}+\frac{9}{2}\cdot\alpha^{5}-\frac{9}{2}\cdot\alpha^{4}-\frac{7}{2}\cdot\alpha^{2}+\frac{3}{2}\cdot\alpha+\frac{1}{2})\cdot t^{2}\\ +(-\frac{3}{2}\cdot\alpha^{6}+2\cdot\alpha^{5}-\frac{3}{2}\cdot\alpha^{4}-2\cdot\alpha^{3}-\frac{3}{2}\cdot\alpha^{2}-\frac{3}{2})\cdot t\\ +\frac{1}{2}\cdot\alpha^{5}-\alpha^{3}+2\cdot\alpha^{2}+\frac{1}{2}\cdot\alpha\end{matrix}\right)\\ \hline\cr\left(\begin{matrix}(-\alpha^{7}-\alpha^{6}-\frac{1}{2}\cdot\alpha^{5}-6\cdot\alpha^{3}-6\cdot\alpha^{2}-\frac{5}{2}\cdot\alpha)\cdot t^{4}\\ +(-\frac{7}{2}\cdot\alpha^{7}-\frac{5}{2}\cdot\alpha^{6}-\frac{1}{2}\cdot\alpha^{5}+\frac{1}{2}\cdot\alpha^{4}-\frac{41}{2}\cdot\alpha^{3}-\frac{29}{2}\cdot\alpha^{2}-\frac{7}{2}\cdot\alpha+\frac{5}{2})\cdot t^{3}\\ +(-5\cdot\alpha^{7}-\frac{5}{2}\cdot\alpha^{6}+\alpha^{4}-29\cdot\alpha^{3}-\frac{29}{2}\cdot\alpha^{2}+6)\cdot t^{2}\\ +(-\frac{7}{2}\cdot\alpha^{7}-\alpha^{6}+\frac{1}{2}\cdot\alpha^{5}+\alpha^{4}-\frac{41}{2}\cdot\alpha^{3}-6\cdot\alpha^{2}+\frac{7}{2}\cdot\alpha+6)\cdot t\\ -\alpha^{7}+\frac{1}{2}\cdot\alpha^{5}+\frac{1}{2}\cdot\alpha^{4}-6\cdot\alpha^{3}+\frac{5}{2}\cdot\alpha+\frac{5}{2}\end{matrix},\right.&\left.\begin{matrix}(\frac{3}{2}\cdot\alpha^{6}+\frac{3}{2}\cdot\alpha^{5}+\frac{1}{2}\cdot\alpha^{4}+\frac{17}{2}\cdot\alpha^{2}+\frac{17}{2}\cdot\alpha+\frac{7}{2})\cdot t^{6}\\ +(5\cdot\alpha^{7}+\frac{15}{2}\cdot\alpha^{6}+5\cdot\alpha^{5}+\frac{3}{2}\cdot\alpha^{4}+29\cdot\alpha^{3}+\frac{87}{2}\cdot\alpha^{2}+29\cdot\alpha+\frac{15}{2})\cdot t^{5}\\ +(18\cdot\alpha^{7}+16\cdot\alpha^{6}+\frac{15}{2}\cdot\alpha^{5}+105\cdot\alpha^{3}+94\cdot\alpha^{2}+\frac{87}{2}\cdot\alpha)\cdot t^{4}\\ +(29\cdot\alpha^{7}+\frac{37}{2}\cdot\alpha^{6}+5\cdot\alpha^{5}-\frac{7}{2}\cdot\alpha^{4}+169\cdot\alpha^{3}+\frac{215}{2}\cdot\alpha^{2}+29\cdot\alpha-\frac{37}{2})\cdot t^{3}\\ +(\frac{51}{2}\cdot\alpha^{7}+\frac{23}{2}\cdot\alpha^{6}-\frac{9}{2}\cdot\alpha^{4}+\frac{297}{2}\cdot\alpha^{3}+\frac{133}{2}\cdot\alpha^{2}-\frac{55}{2})\cdot t^{2}\\ +(12\cdot\alpha^{7}+3\cdot\alpha^{6}-2\cdot\alpha^{5}-3\cdot\alpha^{4}+70\cdot\alpha^{3}+18\cdot\alpha^{2}-12\cdot\alpha-18)\cdot t\\ +\frac{5}{2}\cdot\alpha^{7}-\alpha^{5}-\alpha^{4}+\frac{29}{2}\cdot\alpha^{3}-6\cdot\alpha-5\end{matrix}\right)\\ \hline\cr\left(\begin{matrix}(\frac{1}{2}\cdot\alpha^{7}+\frac{5}{2}\cdot\alpha^{3}-\alpha^{2}-\alpha)\cdot t^{4}\\ +(\frac{1}{2}\cdot\alpha^{7}-\frac{1}{2}\cdot\alpha^{6}-\frac{1}{2}\cdot\alpha^{5}-\frac{1}{2}\cdot\alpha^{4}+\frac{7}{2}\cdot\alpha^{3}-\frac{1}{2}\cdot\alpha^{2}+\frac{1}{2}\cdot\alpha-\frac{1}{2})\cdot t^{3}\\ +(\frac{1}{2}\cdot\alpha^{6}+\alpha^{5}+\alpha^{4}+\frac{1}{2}\cdot\alpha^{2}+\alpha)\cdot t^{2}\\ +(-\frac{1}{2}\cdot\alpha^{7}-\frac{1}{2}\cdot\alpha^{5}-\alpha^{4}-\frac{7}{2}\cdot\alpha^{3}+\alpha^{2}+\frac{1}{2}\cdot\alpha)\cdot t\\ -\frac{1}{2}\cdot\alpha^{7}+\frac{1}{2}\cdot\alpha^{4}-\frac{5}{2}\cdot\alpha^{3}-\alpha+\frac{1}{2}\end{matrix},\right.&\left.\begin{matrix}(\frac{1}{2}\cdot\alpha^{6}+\frac{1}{2}\cdot\alpha^{5}+\frac{1}{2}\cdot\alpha^{4}+\frac{3}{2}\cdot\alpha^{2}-\frac{1}{2}\cdot\alpha-\frac{1}{2})\cdot t^{6}\\ +(-2\cdot\alpha^{5}-3\cdot\alpha^{4}-2\cdot\alpha^{3}+3\cdot\alpha^{2})\cdot t^{5}\\ +(-\frac{3}{2}\cdot\alpha^{7}+3\cdot\alpha^{5}+7\cdot\alpha^{4}-\frac{3}{2}\cdot\alpha^{3}+3)\cdot t^{4}\\ +(2\cdot\alpha^{7}+\frac{1}{2}\cdot\alpha^{6}-2\cdot\alpha^{5}-\frac{15}{2}\cdot\alpha^{4}-\frac{9}{2}\cdot\alpha^{2}-\frac{1}{2})\cdot t^{3}\\ +(-\frac{3}{2}\cdot\alpha^{7}-\frac{5}{2}\cdot\alpha^{6}+\frac{9}{2}\cdot\alpha^{4}+\frac{3}{2}\cdot\alpha^{3}-\frac{7}{2}\cdot\alpha^{2}-\frac{1}{2})\cdot t^{2}\\ +(\alpha^{7}+\frac{3}{2}\cdot\alpha^{6}+\alpha^{5}-\frac{3}{2}\cdot\alpha^{4}+\alpha^{3}+\frac{3}{2}\cdot\alpha^{2}+\alpha-\frac{3}{2})\cdot t\\ -\frac{1}{2}\cdot\alpha^{5}+\alpha^{3}+2\cdot\alpha^{2}-\frac{1}{2}\cdot\alpha\end{matrix}\right)\end{matrix}$
Figure 1. Liftings of the rational points in (11) on $\overline{E}_{1}$ from Section 4.2 to characteristic zero
[…]

### v2

@@
 We denote by $X_{OY}$ the universal cover of $Y_{OY}$.
+Figure 1. The real points of the affine birational model for the Enriques surface sought for by Oguiso and Yu
 Theorem 1.1.
-The Oguiso-Yu K3 surface $X_{OY}$ is the elliptic K3 surface whose weierstrass equation is
+The Oguiso-Yu Enriques surface is birational to the surface with equation
+$\tilde{y}^{2}=s\tilde{x}^{4}+s^{7}-s^{3}.$
+Its universal cover $X_{OY}$ is the elliptic K3 surface whose weierstrass equation is
 $y^{2}=x^{3}+(1-t^{8})x.$
@@
 Acknowledgements
+Special thanks goes to Matthias Schütt for discussions and suggesting the equation for the Enriques surface after reading a first version of this paper.
 The first named author would like to thank Janko Böhm for sharing his insights.
@@
 From the exponential sequence, we thus have an embedding $c_{1}\colon\mathrm{Pic}(X)\to H^{2}(X,\mathbb{Z})$ whose image is the subgroup of integral $(1,1)$-classes $H^{1,1}(X)\cap H^{2}(X,\mathbb{Z})$ by Lefschetz’ $(1,1)$-theorem.
-The smallest primitive sublattice of $H^{2}(X,\mathbb{Z})$ containing the period $H^{2,0}(X)$ is called the transcendental lattice $T(X)$.
+The smallest primitive sublattice of $H^{2}(X,\mathbb{Z})$ whose complexification contains the period $H^{2,0}(X)$, is called the transcendental lattice $T(X)$.
 A K3 surface is not necessarily projective, but if it is, then $T(X)=\mathrm{NS}(X)^{\perp}$ in $H^{2}(X,\mathbb{Z})$.
@@
 which preserves the intersection form and the Hodge structures, i.e.
-$\phi_{\mathbb{C}}(H^{2,0}(Y))=H^{0,2}(X)$.
-The weak Torelli theorem states that $X$ and $Y$ are isometric, if and only if $H^{2}(X,\mathbb{Z})$ and $H^{2}(Y,\mathbb{Z})$ are Hodge-isometric.
-To formulate the strong Torelli theorem, we need one more definition. The Hode-isometry $\phi$ is called effective, if it satisfies one of the following three equivalent conditions:
+$\phi_{\mathbb{C}}(H^{2,0}(Y))=H^{2,0}(X)$.
+The weak Torelli theorem states that $X$ and $Y$ are isomorphic, if and only if $H^{2}(X,\mathbb{Z})$ and $H^{2}(Y,\mathbb{Z})$ are Hodge-isometric.
+To formulate the strong Torelli theorem, we need one more definition. The Hodge-isometry $\phi$ is called effective, if it satisfies one of the following three equivalent conditions:
   1. (1)
@@
 The ample cone is one such chamber.
-The Weyl group $W(X)\subseteq O(\mathrm{NS}(X))$ is generated by the reflections into the hyperplanes $r^{\perp}$.
+The Weyl group $W(X)\subseteq O(\mathrm{NS}(X))$ is generated by the reflections at the hyperplanes $r^{\perp}$.
 It acts simply-transitively on the set of Weyl-chambers.
@@
 A complex Enriques surface $Y$ is a compact complex manifold with $h^{1}(Y,\mathcal{O}_{Y})=0$ and $b_{2}(Y)=10$.
-Its canonical bundle $\omega_{Y}\neq\mathcal{O}_{Y}$, but $\omega_{Y}^{\otimes 2}\cong\mathcal{O}_{Y}$.
+Its canonical bundle is $\omega_{Y}\neq\mathcal{O}_{Y}$, but $\omega_{Y}^{\otimes 2}\cong\mathcal{O}_{Y}$.
 Since the canonical bundle is $2$-torsion, it induces a $2:1$ cover $X\to Y$, which turns out to be the universal covering with $X$ a K3 surface.
@@
 Let $L$ be a quadratic integer lattice and $f\in O(L)$ an isometry. The invariant and co-invariant lattices of $f$ are
-$L^{i}=\{x\in L\mid f(x)=x\}\qquad L_{i}:=(L^{i})^{\perp}.$
+$L^{f}=\{x\in L\mid f(x)=x\}\qquad L_{f}:=(L^{f})^{\perp}.$
 We follow the geometric convention that ADE-root lattices are negative definite.
@@
 Furthermore, we know that $h$ is basepoint free, because, by a direct calculation, the set
-$\{f\in\mathrm{NS}(X)\mid f^{2}=0,h.f=1\}=\emptyset$
+$\{e\in\mathrm{NS}(X)\mid e^{2}=0,h.e=1\}=\emptyset$
 is empty and hence $h$ induces a morphism $\phi:X\to\mathbb{P}^{3}$, cf. [SaintDonat74]. In general, the ample class $h$ can belong to one of the following three types according to [SaintDonat74]:
@@
 The description of the stabilisers and the orbit follows from the proof of Proposition 3.7 as follows: Since $h$ is invariant under $G$, so is the set
-$\{f\in\mathrm{NS}(X)\mid f^{2}=0,h.f=2\}=\{e_{1},e_{2}\}.$
+$\{e\in\mathrm{NS}(X)\mid e^{2}=0,h.e=2\}=\{e_{1},e_{2}\}.$
 Since the invariant lattice is of rank $1$, $G$ must exchange $e_{1}$ and $e_{2}$ and the stabiliser of $e_{i}$ is of order $G/2=128$.
-Recall from [shioda] that the trivial lattice $\mathrm{Triv}(f)$ of an elliptic fibration with fiber $f\in\mathrm{NS}(X)$ is the sublattice of $\mathrm{NS}(X)$ spanned by the zero section and fiber components.
-The Mordell-Weil group is isomorphic to $\mathrm{NS}(X)/\mathrm{Triv}(e_{i})$.
+Recall from [shioda] that the trivial lattice $\mathrm{Triv}(e)$ of an elliptic fibration with fiber $e\in\mathrm{NS}(X)$ is the sublattice of $\mathrm{NS}(X)$ spanned by the zero section and fiber components.
+The Mordell-Weil group is isomorphic to $\mathrm{NS}(X)/\mathrm{Triv}(e)$.
 To compute the trivial lattice, one chooses any $s\in\mathrm{NS}(X)$ with $s^{2}=-2$ and $s.f=1$.
-Let $R$ be the root sublattice of $\langle f,s\rangle^{\perp}\subseteq\mathrm{NS}(X)$.
-Then the trivial lattice is isomorphic to $\mathbb{Z}\langle s,f\rangle\oplus R$.
+Let $R$ be the root sublattice of $\langle e,s\rangle^{\perp}\subseteq\mathrm{NS}(X)$.
+Then the trivial lattice is isomorphic to $\mathbb{Z}\langle s,e\rangle\oplus R$.
 The irreducible root sublattices of $R$ give the ADE types of the singular fibers.
@@
 ∎
-It then follows from Proposition 3.2 that $X_{OY}\cong X\cong X_{h}$ and from Proposition 3.6 that $X_{h}/j=:Y$ is isomorphic to the Oguiso-Yu Enriques surface $Y_{OY}$ of minimum entropy.
+It then follows from Proposition 3.2 that $X_{OY}\cong X\cong X_{h}$ and from Proposition 3.6 that $X_{h}/j=:Y$ is isomorphic to the Oguiso-Yu Enriques surface $Y_{OY}$ with an automorphism of minimum entropy.
 4.2. Search for sections
@@
 The associated elliptically fibered K3 surface $X_{e_{1}}$ is defined over $\mathbb{Q}$ and its trivial lattice over $\mathbb{Q}(\zeta_{8})$.
-However, in order to realize all algebraic cycles for the Néron-Severi lattice $\mathrm{NS}(X_{e_{1}})$, we still need the extension $K/\mathbb{Q}(\zeta_{8})$ of degree two generated by the algebraic element $\alpha$ with minimal polynomial given in Equation 6.
+However, in order to realize all algebraic cycles for the Néron-Severi lattice $\mathrm{NS}(X_{e_{1}})$, we still need the extension $K/\mathbb{Q}(\zeta_{8})$ of degree two, generated by the algebraic element $\alpha$ with minimal polynomial over $\mathbb{Q}$ given in Equation 6.
+Proposition 4.4.
+The surface
+(9) $\tilde{y}^{2}=s\tilde{x}^{4}+s^{7}-s^{3}$
+is birational to the Oguiso-Yu Enriques surface $Y_{OY}$.
+Before we go into the proof, we review how to obtain an elliptic fibration of the covering K3 surface greater generality: Let $\widetilde{\pi}\colon Y\to\mathbb{P}^{1}$ an elliptic fibration of an Enriques surface with double fibers over $P_{1},P_{2}\in\mathbb{P}^{1}$, and let $\kappa\colon X\to Y$ be the universal cover.
+Suppose that $\widetilde{\pi}$ admits a rational bisection.
+According to [cdl:EnriquesI, Proposition 4.10.10] the Stein-factorization of $\widetilde{\pi}\circ\kappa\colon X\to Y\to\mathbb{P}^{1}$ is given by an elliptic fibration $\pi\colon X_{OY}\to\mathbb{P}^{1}$ followed by a $2:1$ cover $\mathbb{P}^{1}\to\mathbb{P}^{1}$ ramified over $P_{1},P_{2}$.
+Proof of Proposition 4.4.
+We set $t^{2}=s$, $\tilde{x}=tx_{1}$, $\tilde{y}=t^{3}y_{1}$ in equation (9) and obtain $y_{1}^{2}=x_{1}^{4}+t^{8}-1$.
+Using the rational point $(x_{1},y_{1})=(1,t^{4})$ we can transform the equation to a short Weierstrass form which one checks to be isomorphic to the one of eq.
+(8).
+Putting everything together, we obtain obtain the universal covering morphism $\kappa_{OY}\colon X_{OY}\to Y_{OY}$ by means of the following rational expressions in the Weierstrass coordinates of $X_{OY}$:
+$\begin{pmatrix}\tilde{x}\\ \tilde{y}\\ s\end{pmatrix}\mapsto\begin{pmatrix}\frac{-x\cdot t^{9}+x\cdot t+\frac{1}{2}\cdot\sqrt{2}\cdot y\cdot t^{5}-t^{9}+t}{x+\frac{1}{2}\cdot\sqrt{2}\cdot y\cdot t^{4}+t^{8}-1}\\ \frac{-x^{3}\cdot t^{15}-3\cdot x^{2}\cdot t^{15}+3\cdot x^{2}\cdot t^{7}+\frac{1}{2}\cdot y^{2}\cdot t^{15}+\sqrt{2}\cdot y\cdot t^{19}-3\cdot\sqrt{2}\cdot y\cdot t^{11}+2\cdot\sqrt{2}\cdot y\cdot t^{3}+t^{23}-2\cdot t^{15}+t^{7}}{x^{2}+\sqrt{2}\cdot x\cdot y\cdot t^{4}+2\cdot x\cdot t^{8}-2\cdot x+\frac{1}{2}\cdot y^{2}\cdot t^{8}+\sqrt{2}\cdot y\cdot t^{12}-\sqrt{2}\cdot y\cdot t^{4}+t^{16}-2\cdot t^{8}+1}\\ t^{2}\end{pmatrix}$
+As ultimate check we confirmed that $\kappa_{OY}=\kappa_{OY}\circ\iota_{OY}$. These computations are carried out in the file
+enriques_projection.jl. ∎
 4.4. Specialisation to positive characteristic